### v2

# Sample Complexity of Diffusion Model Training Without Empirical Risk Minimizer Access

Abstract
Diffusion models have demonstrated state-of-the-art performance across vision, language, and scientific domains. Despite their empirical success, prior theoretical analyses of the sample complexity suffer from poor scaling with input data dimension or rely on unrealistic assumptions such as access to exact empirical risk minimizers. In this work, we provide a principled analysis of score estimation, establishing a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$. Our approach leverages a structured decomposition of the score estimation error into statistical, approximation, and optimization errors, enabling us to eliminate the exponential dependence on neural network parameters that arises in prior analyses. It is the first such result which achieves sample complexity bounds without assuming access to the empirical risk minimizer of score function estimation loss.
1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks [38, 4], text generation [25], sequential data modeling [37], reinforcement learning and control [42], and life sciences [19, 30]. For a more comprehensive exposition of applications, we refer readers to survey paper [9].
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. The iteration complexity or convergence has been studied in [23, 5, 24, 18, 11, 26, 27], while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function. We note that the [40, 39, 32, 8] have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, [15] did an analysis as a baseline for the approach in [7], and showed that this approach can have polynomial dependence in $\epsilon$, but have exponential dependence on neural network parameters, making the result less important for large neural networks. The key previous work without an exponential dependence on neural network or dimension is [15] which obtains $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity [^1] We note that [15] claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix B. . However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss a significant restriction that was explicitly highlighted as an open problem in [15]. In this paper, we address this open problem and overcome this limitation by establishing a sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-6})$ without assuming access to the ERM of the score estimation loss. This represents a key step toward bridging the gap between the theory and practice of diffusion models. Specifically, we address the following fundamental question:
How many samples are required for a sufficiently expressive neural network to estimate the score function well enough to generate high-quality samples using a DDPM-type algorithm, without assuming access to the empirical minimizer of the score function estimation loss?
In this paper, we provide a refined analysis that yields a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$, while avoiding the exponential dependence on data dimension and neural network parameters as in [15]. This is achieved through a decomposition of the total variation error into distinct terms associated with forward process approximation, discretization of the reverse-time process, and score estimation. Using our novel analysis of the score estimation error, we obtain the sample complexity bounds without exponential dependence on data dimension. Our principled analysis accounts for both statistical and optimization errors while not assuming the access to the empirical risk minimizer of the score estimation loss, and achieving above mentioned sample complexity bounds.
Reference
Sample
Complexity
Empirical Risk
Minimizer Assumption
Exponential Dependence
on Data Dimension
[40] $\tilde{O}\left(\epsilon^{-d}\right)$ Yes Yes
[39] $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes Yes
[32] $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes Yes
[8] $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes Yes
[15]1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes No
This work $\tilde{O}(\epsilon^{-6})$ No No
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. The additional $\tilde{\mathcal{O}}(\epsilon^{-1})$ factor in our sample complexity bound arises from relaxing the ERM assumption on the score estimation loss. This should not be interpreted as a weaker result, but rather as a reflection of the inherent trade-off between making the ERM assumption and the resulting sample complexity bounds. Here $d$ represents the data dimension. For further details on how the sample complexity bounds are derived for [15], see Appendix B.
The score estimation is bounded to account for approximation error due to the limited expressiveness of the function class, statistical error due to finite sample size, and the optimization error due to not reaching global minimum during training. We note that the dimension-dependent term is the approximation error, which can be bounded for over-parametrized neural networks. This work provides the first analysis in this area that explicitly quantifies the optimization error, introducing a novel recursive analysis of the error at each stochastic gradient descent (SGD) step. This approach captures the effect of using a finite number of SGD iterations to estimate the score function.
Furthermore, this work introduces a novel analysis of the statistical error that avoids the exponential dependence on neural network parameters observed in [7], and does not rely on the quantile-based formulation used in [15]. This is achieved by directly bounding the gap between the expected and empirical loss over a finite function class, thereby sidestepping the need for bounds that scale poorly with network size. A key step in our analysis leverages the quadratic growth property implied by the Polyak-Łojasiewicz (PL) condition assumed in Assumption 2, which governs the score estimation error. This condition plays a critical role in controlling the sample complexity. While stronger growth conditions (e.g., lower Lipshitz) could in principle yield a sample complexity of $\widetilde{\mathcal{O}}(\epsilon^{-3})$, standard neural networks typically satisfy only the PL condition. Moreover, recent work [15] assumes access to the empirical risk minimizer (ERM) of the score estimation loss an assumption that is unrealistic in practice. Under this assumption, they achieve a sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-5})$. In contrast, by removing this impractical assumption and instead relying on the Polyak-Łojasiewicz (PL) condition, our analysis yields a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$. This highlights a trade-off between assuming access to an ERM and the resulting sample complexity: relaxing the ERM assumption comes at the cost of a slightly worse (yet more realistic) sample bound.
The main contributions of our work are summarized as:
  • Finite time sample complexity bounds. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$ for score-based diffusion models, and do not have exponential dependence on the data dimension or neural network parameters. Our analysis avoids the unrealistic assumptions used in prior work such as access to empirical loss minimizer.
  • Principled error decomposition. We propose a structured decomposition of the score estimation error into approximation, statistical, and optimization components, enabling the characterization of how each factor contributes to sample complexity.
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in [10] showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works [23, 5, 24, 18, 11, 26, 27]. In particular, [5, 23] establish iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates [22, 27, 11]. Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in [24, 18, 26]. In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the error introduced by neural network approximation, data sampling, and optimization are also accounted.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in [8, 40, 39, 32]. However, these bounds are exponential in the data dimension. Further, authors of [15] use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
2 Preliminaries and Problem Formulation
We begin by outlining the theoretical basis of score-based diffusion models. In particular, we adopt the continuous-time stochastic differential equation (SDE) framework, which provides a principled basis for modeling the generative process. We then outline its practical discretization and formally define our problem.
Score-based generative models enable sampling from a complex distribution $p_{0}$ by learning to reverse a noise-adding diffusion process. This approach introduces a continuous-time stochastic process that incrementally perturbs the data distribution into a tractable distribution (typically Gaussian), and then seeks to reverse that transformation.
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process [33], defined by the following SDE
$dx_{t}=-x_{t}dt+\sqrt{2}dB_{t},\quad x_{0}\sim p_{0},x\subset\mathbb{R}^{d},$ (1)
where $B_{t}$ denotes standard Brownian motion. The solution of this SDE in the closed form is given by
$x_{t}\sim e^{-t}x_{0}+\mathcal{N}(0,\sigma_{t}^{2}I),\quad\text{with }\sigma_{t}^{2}=1-e^{-2t}.$ (2)
As $t\to\infty$, the process converges to the stationary distribution $\mathcal{N}(0,I)$. Let $p_{t}$ denote the marginal distribution of $x_{t}$. This defines a continuous-time smoothing of the data distribution, where $p_{t}$ becomes increasingly Gaussian over time.
This is typically achieved using stochastic time-reversal theory [3], which yields a corresponding reverse-time SDE as follows.
$dx_{T-t}=\left(x_{T-t}+2\nabla\log p_{T-t}(x_{T-t})\right)dt+\sqrt{2}\,dB_{t},$ (3)
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions.
Assuming the second moment of $p_{0}$ is bounded, i.e., $m_{2}^{2}:=\mathbb{E}_{x\sim p_{0}}[\|x\|^{2}]$, it is sufficient to choose $T=\mathcal{O}(\log(m_{2}^{2}/\epsilon))$ to ensure that the reverse process generates samples within $\epsilon$-total variation distance of $p_{0}$ [6].
In practice, the continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework [16], where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix C. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
Problem formulation: Let the score function be approximated using a parameterized family of functions $\mathcal{F}_{\Theta}=\{s_{\theta}:\theta\in\Theta\}$, where each $s_{\theta}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ is typically represented by a neural network of depth $D$ and width $W$ with smooth activation functions. Given $n$ i.i.d. samples $\{x_{i}\}_{i=1}^{n}$ from the data distribution $p_{t}$, the score network is trained by minimizing the following time-indexed loss:
$\mathcal{L}_{t}(\theta):=\mathbb{E}_{x_{t}\sim p_{t}}\left[\|s_{\theta}(x_{t},t)-\nabla\log p_{t}(x_{t})\|^{2}\right].$ (4)
Let $p$ be the original data distribution. We additionally define the following probability distributions:
$p_{t_{0}}\coloneqq$ Distribution obtained after backward diffusion till time $t_{0}$ steps starting form $p_{T}$
$p^{dis}_{t_{0}}\coloneqq$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
at discretized time steps
$\tilde{p}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
at discretized time steps using the estimated score functions
$\hat{p}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $\mathcal{N}(0,I)$
at discretized time steps using the estimated score functions
where $t_{0}$ denotes the early stopping time.
Objective. Our goal is to quantify how well the learned generative model $\hat{p}_{t_{0}}$ approximates the true data distribution $p$ in terms of total variation (TV) distance. Specifically, we aim to show the number of samples needed so that with high probability, the TV distance $\mathrm{TV}(p,\hat{p}_{t_{0}})$ is bounded by $\mathcal{O}(\epsilon)$, where $\epsilon$ is the $L^{2}$ estimation error of the score function.
3 Sample Complexity of Diffusion Models
In this section, we derive explicit sample complexity bounds for diffusion-based generative models. By leveraging tools from stochastic optimization and statistical learning theory, we provide bounds on the number of data samples required to accurately estimate the time-dependent score function $s_{t}(x):=\nabla\log p_{t}(x)$ across the forward diffusion process. Note that accurate score estimation is critical for ensuring high-quality generation while sampling through the reverse-time SDE.
We first state the assumptions required throughout this work. We begin with a commonly used smoothness condition [17, 21] on the score function, which ensures well-behaved gradients and facilitates convergence guarantees for stochastic first-order methods.
Assumption 1 (Smoothness and bounded gradient variance of the score loss.).
For all $t\in[0,T]$, the population loss $\mathcal{L}_{t}(\theta)$ is $\kappa$-smooth with respect to the parameters $\theta$, i.e., for all $\theta,\theta^{\prime}\in\Theta$
$\|\nabla\mathcal{L}_{t}(\theta)-\nabla\mathcal{L}_{t}(\theta^{\prime})\|\leq\kappa\|\theta-\theta^{\prime}\|.$ (5)
We assume that the estimators of the gradients ${\nabla}\mathcal{L}_{t}(\theta)$ have bounded variance.
[…]
The bounded gradient variance assumption is a standard assumption used in SGD convergence results such as [20] and [1]. This assumption holds for a broad class of neural networks under mild conditions, such as Lipschitz-continuous activations (e.g., GELU), bounded inputs, and standard initialization schemes [2, 12]. The smoothness condition ensures that the gradients of the loss function do not change abruptly, which is crucial for stable updates during optimization. This stability facilitates a controlled optimization dynamics when using first-order methods like SGD or Adam, ensuring that the learning process proceeds in a well-behaved manner. We emphasize that this smoothness assumption and bounded gradient assumptions are not present in prior works such as [15] and [7], which evaluate the sample complexity of diffusion models under the assumption of access to the empirical risk minimizer, rather than SGD based optimization.
Assumption 2 (Polyak - Łojasiewicz (PL) condition.).
The loss $\mathcal{L}_{t}(\theta)$ for all $t\in[0,T]$ satisfies the Polyak–Łojasiewicz condition, i.e., there exists a constant $\mu_{t}>0$ such that
$\frac{1}{2}\|\nabla\mathcal{L}_{t}(\theta)\|^{2}\geq\mu_{t}\left(\mathcal{L}_{t}(\theta)-\mathcal{L}_{t}(\theta^{*})\right),\quad\forall~{}\theta\in\Theta,$ (7)
where $\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}_{t}(\theta)$ denotes the global minimizer of the population loss.
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses [29]. Prior works such as [15] and [7] implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in [15] and the definition of $\hat{f}$ in Theorem 13 of [7]). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.
In contrast, the PL condition allows us to derive sample complexity bounds under realistic optimization dynamics, without requiring exact ERM solutions. To our knowledge, this is the first theoretical analysis of score-based generative models that explicitly accounts for inexact optimization, addressing a key gap in existing literature. Additionally, we establish convergence guarantees with both constant and decreasing step sizes.
Assumption 3 (Sub-Gaussian Data Distribution.).
The data distribution $p_{0}$ of the data variable $x_{0}$ has an absolutely continuous CDF, is supported on a continuous set $\Gamma\in\mathbb{R}^{d}$, and there exists a constant $0<C_{1}<\infty$ such that for every $t\geq 0$ we have $P(|x_{0}|\geq t)\leq 2\cdot{\exp}^{-\frac{t^{2}}{C^{2}_{1}}}$.
[10, 32] assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to be sub-Gaussian, making our results applicable to a significantly broader class of distributions.
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function [28]. Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models [34, 29]. Under these assumptions, we now characterize the sample complexity of score-based learning in the following.
Main Result: To estimate the score function $s_{t}(x)$, modern diffusion models train a neural network to predict the noise $\epsilon_{\theta}(x_{t},t)$ injected in the forward process. The training objective, as used in DDPM [16], corresponds to minimizing
$\mathcal{L}(\theta)=\mathbb{E}_{x,t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (8)
where $x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noisy observation at time $t$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ is defined by the noise schedule $\{\beta_{t}\}_{t=1}^{T}$. This objective aligns with the score-matching loss under a reparameterized formulation. Algorithm 1 (see Appendix C) outlines the standard DDPM training and sampling pipeline. The final generative quality depends on the fidelity of the learned score estimate across time steps.
The following theorem provides an upper bound on the total variation distance between the true data distribution $p$ and the generated distribution $\hat{p}_{t_{0}}$.
Theorem 1 (Total Variation Distance Bound).
Let $p(x)$ denote the true data distribution, and $\hat{p}_{t_{0}}(x)$ be the distribution generated by a discretized diffusion model with time steps $\{t_{k}\}_{k=0}^{K}$, reverse-time score estimates $\hat{s}_{T-t_{k}}(x)$, and stopping at time $t_{0}$. Let $d$ be the data dimension, and $n_{k}$ be the number of samples for score estimation at time step $t_{k}$.
Assume the loss function $\mathcal{L}_{t}(\theta)$ satisfies Assumptions 1 and 2, the data distribution satisfies Assumption 3, and the learning rate satisfies $0\leq\eta\leq\frac{1}{\kappa}$. Further assume
$n_{k}=\mathcal{O}\left(d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-6}}{\sigma_{k}^{-4}}\right)\right),$ (9)
and that the neural network has width $W$ and depth $D$ such that $WD=\mathcal{O}\left((\epsilon/\sigma_{k})^{-\frac{d}{6}}\right)$.
Then, with probability at least $1-\delta$, the total variation distance between the true distribution $p$ and $\hat{p}_{t_{0}}$ satisfies
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\left(\epsilon\cdot{T}\right)+\epsilon$ (10)
Furthermore, by setting $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, $t_{0}=\tilde{\mathcal{O}}(\epsilon^{2})$ and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\tilde{\mathcal{O}}(\epsilon),$ (11)
with probability at least $1-\delta$.
Theorem 1 establishes that the total variation distance between the true data distribution and the diffusion model’s output can be made arbitrarily small specifically, $\tilde{\mathcal{O}}(\epsilon)$ by properly scaling model capacity and algorithmic parameters. To the best of our knowledge, these are the only known sample complexity bounds for score-based diffusion models, improving upon the prior results as discussed in the introduction without assuming access to empirical risk minimizer for the score estimation loss. Our result highlight an inherent trade-off between the sample complexity bounds and the assumption of access to the empirical risk minimizer.
Proof of Theorem 1. Recall that $\hat{p}_{t_{0}}$ is derived via score-based sampling, so using the triangle inequality repeatedly to decompose the TV distance between the true distribution $p$ and $\hat{p}_{t_{0}}$
Using the triangle property of $\mathrm{TV}$ we write
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\mathrm{TV}(p,p_{t_{0}})+\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$ (12)
$\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$ (13)
The bound on $\mathrm{TV}(p,p_{t_{0}})$ follows from Lemma D.5 of [13].
Now lemma B.2 of [15] states that if $P\left(\sum_{k=0}^{K}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}.(t_{k+1}-t_{k})\geq\epsilon^{2}\right)\leq\beta$, where $\hat{s}_{t_{k}}(x,t_{k})$ is the estimate of the score ${\nabla}{\log}p_{t_{k}}(x)$, then we have
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\beta+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (14)
Let ${\beta(n)}$ be such that
$\sum_{k=t_{o}}^{K}\mathbb{E}_{x\sim p_{t_{k}}}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}(t_{k}-t_{k-1})\leq{\beta(n)}$ (15)
From Theorem 2 from [31] we have that
$P(\sum_{k=t_{o}}^{K}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}\geq\epsilon^{2})\leq\frac{1}{\epsilon^{2}}\cdot\beta(n)$ (16)
Here $n$ denotes the number of samples used to estimate the loss functions $\hat{s}_{t_{k}}(x,t_{k})$ Thus, we obtain that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\delta+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (17)
$\leq\beta(n).\frac{1}{\epsilon}+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (18)
The bounds on $\mathrm{TV}(p^{\prime}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ and $\mathrm{TV}(p_{t_{0}},p^{\prime}_{t_{0}})$ follow from Lemma B.4 of [15] and Proposition 4 of [5], respectively to get
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\exp(-T)+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)+\beta(n)\cdot\frac{1}{\epsilon^{2}}+\epsilon$ (19)
Note that we have used results from [15], [5] and [13]. [15] and [5] assume a bounded second moment for the data distribution and [13] assumes sub-Gaussian structure for the data distribution. Both of these assumptions are together given in Assumption 3. Putting the expressions for $\mathrm{TV}(p^{\prime}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$, and $\mathrm{TV}(p_{t_{0}},p^{\prime}_{t_{0}})$ in Equation (18), we get
A key point here is that in [15], the term $\sum_{k=0}^{K}\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$ was required to be bounded by $\epsilon^{2}$ with high probability in order for the upper bound on the total variation distance between the learned and true distributions to hold. In contrast, we relax this requirement by converting the probabilistic bound into a bound on the expectation of the same quantity. Specifically, we show that it suffices for the expectation of $\sum_{k=0}^{K}\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$ to be upper bounded by $\beta(n)$ (See Remark 1 for detail). This shift from high-probability bounds to an expectation-based analysis is a crucial aspect of our approach. In the following, we analyze how $\beta(n)$ depends on the number of samples used to estimate the score function $\hat{s}_{t_{k}}(x)$.
In order to obtain an expression for $\beta(n)$, we denote $A(k)$ as
$A(k)\coloneqq E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}dx$ (20)
Now, for each time step $k$, we decompose the total score estimation error, denoted by $A(k)$, into three primary components: approximation error, statistical error, and optimization error. Each of these error corresponds to a distinct aspect of learning the reverse-time score function in a diffusion model as described below.
$\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right]$ $\leq 4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|s^{a}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{approx}}_{k}}$
$\quad+4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{stat}}_{k}}$
$\quad+4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|\hat{s}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{opt}}_{k}},$ (21)
where, we define the parameters
$\theta^{a}_{k}$ $=\arg\min_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t}(x,t_{k})\|^{2}\right],$ (22)
$\theta^{b}_{k}$ $=\arg\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i},t_{k})\right\|^{2}$ (23)
and denote $s^{a}_{k}$ and $s^{b}_{k}$ as the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$ respectively. Approximation error $\mathcal{E}^{\mathrm{approx}}_{k}$ captures the error due to the limited expressiveness of the function class $\{s_{\theta}\}_{\theta\in\Theta}$. The statistical error $\mathcal{E}^{\mathrm{stat}}_{k}$ is the error from using a finite sample size. Finally, the optimization error $\mathcal{E}^{\mathrm{opt}}_{k}$ is due to not reaching the global minimum during training.
One of our key contributions lies in rigorously bounding each of these error components and showing how their interplay governs the overall generative error. In particular, we derive novel bounds that explicitly capture the dependencies on sample size, neural network capacity, and optimization parameters, without any assumption on the access to empirical risk minimizer of the score estimation loss. We formalize these results in the following lemmas. Detailed proofs are deferred to Appendices D.1, D.2, and D.3, respectively.
Lemma 1 (Approximation Error).
Let $W$ and $D$ denote the width and depth of the neural network architecture, respectively, and let $d$ represent the input data dimension. Then, under Assumption 3 for all $k\in[{0},K]$, we have
$\mathcal{E}^{\mathrm{approx}}_{k}\leq\mathcal{O}\left(\left(WD\right)^{-\frac{2}{d}}\right).$ (24)
For this lemma we use the conditional normality of the score function conditioned on the initial data observation to account for the fact that the score function can be unbounded. We combine that with approximation upper bounds from [36, 41] to obtain our result. The specific form of decomposition of the score estimation error allows this term to be separate from the error terms depending on the number of samples used for score estimation. This term also accounts for the error due to the limited approximation power of a class of neural networks, an important factor that was not considered in previous analysis.
Lemma 2 (Statistical Error).
Let $n_{k}$ denote the number of samples used to estimate the score function at each diffusion step $t_{k}$. Then, under Assumptions 3, for all $k\in[{0},K]$, with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)$ (25)
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Our novel analysis here allows us to avoid the exponential dependence on the neural networks parameters in the sample complexity. This is achieved without resorting to the quantile formulation and assuming access to the empirical risk minimizer of the score estimation loss which leads to a sample complexity of $\widetilde{\mathcal{O}}(\epsilon^{-5})$ in [15].
Lemma 3 (Optimization Error).
Let $n_{k}$ be the number of samples used to estimate the score function at diffusion step ${t_{k}}$. If the learning rate $0\leq\eta\leq\frac{1}{\kappa}$, then under Assumptions 1, 2 and 3, for all $k\in[0,K]$, the optimization error due to imperfect minimization of the training loss satisfies with probability at least $1-\delta$
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right).$ (26)
We leverage assumptions 1, 2, and 3 alongside our unique analysis to derive an upper bound on $\mathcal{E}_{\mathrm{opt}}$. The key here is our recursive analysis of the error at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function. This is the first analysis of diffusion models to explicitly account for this error, whereas prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly.
Combining the decomposition in Equation (21) along with Lemmas 13, we obtain the following bound on $A(k)$ (Equation (20)) with probability at least $1-\delta$
$A(k)$ $\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left(d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right)+\mathcal{O}\left((WD)^{-\frac{2}{d}}\right)$
$\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left((WD)^{-\frac{2}{d}}\right),$ (27)
where in the second inequality we combine the first two term appropriately. Setting the network size $WD=\mathcal{O}\left((\epsilon/\sigma_{k})^{-\frac{d}{6}}\right)$ and the sample size
$n_{k}=\mathcal{O}\left(d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-6}}{\sigma_{k}^{-4}}\right)\right),$ (28)
we ensure that $A(k)\leq\frac{\epsilon^{3}}{\sigma_{k}^{2}}=\frac{\epsilon^{3}}{1-e^{-2(T-t_{k})}}$ for all $k\in\{0,\dots,K-1\}$. Summing over all time steps, we obtain with probability at least $1-\delta$
[…]
Note that the term $\left(\log^{2}\left(\frac{4K}{\delta}\right)\right)$ appears in the upper bound for $n_{k}$ in Equation (28) since we have to take a union bound for Lemma 2 and Lemma 3 and then take a union bound over $K$ discretization steps. Substituting this bound into Equation (20), and subsequently into the decomposition in Equation (14) we get that $\beta(n)\leq\epsilon^{3}T$. Substituting this bound in Equation (19) we obtain that with probability at least $1-\delta$.
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\epsilon\cdot{T}+{\epsilon}$ (30)
Finally, by choosing $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, $t_{0}=\tilde{\mathcal{O}}(\epsilon^{2})$ and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we conclude that with probability at least $1-\delta$
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\tilde{\mathcal{O}}(\epsilon),$ (31)
completing the proof of Theorem 1. ∎
Remark 1.
The key distinction between our approach and [15] lies in the error aggregation strategy. In [15], each term $\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}$ in the summation (Equation (15)) must satisfy $\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}\leq\epsilon^{2}$ with probability at least $1-\frac{\beta}{K}$ for all $k\in[0,K]$, necessitating a union bound over $K$ terms (see Lemma C.1 in [15]). This introduces an extraneous factor of $K$ in their sample complexity (evident in Theorem C.2 of [15], detailed in Appendix B). In contrast, we bound the expected cumulative error $\sum_{k=0}^{K}\mathbb{E}_{x\sim p_{t_{k}}}\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$.
Error Decomposition: The expectation-centric analysis enables principled decomposition of the score estimation error into approximation, statistical, and optimization components (Equation (21)). This framework explicitly quantifies errors arising from imperfect optimization, a feature absent in previous works, including [15], which assume access to exact empirical risk minimizers.
In summary, our work provides a principled decomposition of the errors in score-based generative models, highlighting how each component contributes to the overall sample complexity. This leads to the first finite sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$ for diffusion models without assuming access to the empirical minimizer of the score estimation function. The sample complexity bound also avoids the exponential dependence on the data dimension or the neural network parameters.
4 Conclusion and Future Work
In this work, we investigate the sample complexity of training diffusion models via score estimation using neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$, which, to our knowledge, is the first such result that does not assume access to an empirical risk minimizer of the score estimation loss. Notably, our bound does not depend exponentially on the number of neural network parameters. For comparison, the best-known existing result achieves a bound of $\widetilde{\mathcal{O}}(\epsilon^{-5})$, but it crucially assumes access to an ERM. All prior results establishing sample complexity bounds for diffusion models have made this assumption. Our contribution is the first to establish a sample complexity bound for diffusion models under the more realistic setting where exact access to empirical risk minimizer of the score estimation loss is not available.
While our analysis focuses on continuous distributions, extending these guarantees to discrete settings remains an important direction for future work. A promising direction would be to extend our results to conditional diffusion models such as classifier free and classifier guided diffusion.
Appendix A Appendix
This appendix contains additional technical material, including full proofs of our main results, detailed explanations of key lemmas, algorithmic descriptions, and further discussion to support and extend the main text.
Appendix B Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in [15], and show how combining their results with those of [7] leads to an alternative bound.
B.1 Sample Complexity of [15]
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in [15]. A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of [15] is as follows. We re-iterate this Theorem here.
Theorem C.2 [15]. Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
$\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|s_{\theta}(x)-s_{t_{k}}(x)\right\|_{2}^{2}\right]\leq\frac{\delta\cdot\epsilon^{3}}{CK^{2}\sigma_{T-t_{k}}^{2}}\cdot\frac{1}{\log\frac{d+m_{2}+1/m_{2}}{\gamma}}$ (32)
for sufficiently large constant $C$, then consider the score functions trained from
$m>\widetilde{\mathcal{O}}\left(\frac{K(d+\log\frac{1}{\delta})\cdot PD}{\epsilon^{3}}\cdot\log\left(\frac{\max(m_{2},1)\cdot\Theta}{\delta}\right)\cdot\log\left(\frac{m_{2}+1/m_{2}}{\gamma}\right)\right).$ (33)
i.i.d. samples of $q$, with $1-\delta$ probability, DDPM can sample from a distribution $\epsilon$-close in TV to a distribution $\gamma m_{2}$-close in 2-Wasserstein to $q$ in $N$ steps.
Note that Lemma B.6 of [15] states that
$\mathrm{TV}(p,\hat{p})\leq{\delta}+\mathcal{O}\left(\frac{1}{\sqrt{N}}\right)+{\epsilon}\cdot{\sqrt{T}}+\mathcal{O}(\exp^{-T})$ (34)
Here $p$ and $\hat{p}$ are the true and learned data distributions, respectively.
In order to achieve $\mathrm{TV}(p,\hat{p})\leq\epsilon$ we have to set $\delta=\epsilon$. This would imply $N=\mathcal{O}(\epsilon^{-2})$ and $T=\mathcal{O}\left(\frac{1}{\epsilon}\right)$. Putiing this value of $N$ in Equation (33) we obtain that for
$m>\widetilde{O}\left(\frac{(d+\log(1/\delta))\cdot PD}{\epsilon^{5}}\cdot\log\left(\frac{\Theta}{\epsilon}\right)\right).$ (35)
we have with probability at least $1-\epsilon$
$\mathrm{TV}(p,\hat{p})\leq\epsilon$ (36)
This reveals a discrepancy between the reported sample complexity and the actual bound derived above, highlighting that the true complexity is significantly higher than what was originally reported.
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike [15], our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
B.2 A note on the sample complexity result in [7]
The authors of [15] cite the intermediate results of [7] which obtained a sample complexity result of $\tilde{O}(d^{\frac{5}{2}}\cdot W^{D/2}\cdot\epsilon^{-2})$ for a single discretization point. In Section 4 of [15] it is explained that simply obtaining an upper bound on the L2 error at one discretization point is not sufficient to obtain a TV bound between the learned and true probability distributions with probability. This is formally stated in Lemmas 4.1 and 4.2 where it is proved that even when the true distribution is a simple Gaussian, there is a lower bound on the L2 error between the true and estimated score function. Thus the sample complexity of $\tilde{O}(d^{\frac{5}{2}}\cdot W^{D/2}\cdot\epsilon^{-2})$ proved in [7] does not ensure convergence in TV between the true and estimated probability distributions. Additionally, the sample complexity result in [7] is exponential in the depth of of the score estimation neural network, something which is not the case in the sample complexity result we have achieved in this work.
Appendix C Score Estimation Algorithm
In this section, we provide a detailed description of the algorithm used for estimating the score function in diffusion models.
Algorithm 1 Denoising Diffusion Probabilistic Model (DDPM)
1:  Input: Dataset $\mathcal{D}$, timesteps $T$, stop time $t_{\text{stop}}$, schedule $\{\beta_{t}\}_{t=1}^{T}$, network $\epsilon_{\theta}$, learning rate $\eta$, iterations $K$
2:  Precompute: $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ Training (Score Estimation)
3:  for $k=1$ to $n$ do
4:     Sample $x_{i}\sim\mathcal{D}$, $t_{i}\sim\mathrm{Uniform}([1,T])$, $\epsilon_{i}\sim\mathcal{N}(0,I)$ for $i=1,\dots,B$
5:     $x_{t_{i}}=\sqrt{\bar{\alpha}_{t_{i}}}x_{i}+\sqrt{1-\bar{\alpha}_{t_{i}}}\epsilon_{i}$
6:     Compute loss: $\hat{L}(\theta)=\|\epsilon_{i}-\epsilon_{\theta}(x_{t_{i}},t_{i})\|^{2}$
7:     Update $\theta\leftarrow\theta-\eta_{k}\cdot\nabla_{\theta}\hat{L}(\theta)$
8:  end for Sampling
9:  Sample $x_{T}\sim\mathcal{N}(0,I)$
10:  for $t=T$ down to $t_{\text{stop}}+1$ do
11:     $z\sim\mathcal{N}(0,I)$ if $t>1$ else $z=0$
12:     $\hat{\epsilon}=\epsilon_{\theta}(x_{t},t)$
13:     $\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}\right)$
14:     $x_{t-1}=\tilde{\mu}_{t}+\sqrt{\beta_{t}}\cdot z$
15:  end for
16:  Return $x_{t_{\text{stop}}}$
Appendix D Proofs of Intermediate Lemmas
In this section, we present the proofs of intermediate lemmas used to bound the approximation error, statistical error, and optimization error in our analysis.
D.1 Bounding the Approximation Error
Recall Lemma 1 (Approximation Error.) Let $W$ and $D$ denote the width and depth of the neural network architecture, respectively, and let $d$ represent the input data dimension. Then, under Assumption 3 for all $k\in[{0},K]$ we have
$\mathcal{E}^{\mathrm{approx}}_{t}\leq\mathcal{O}\left(\left(WD\right)^{-\frac{2}{d}}\right).$ (37)
Proof.
Recall that the approximation error $\mathcal{E}^{\mathrm{approx}}_{t}$ captures the inherent expressiveness limitations of the function class parameterized by the neural network. Even with infinite data and perfect optimization, the network may not exactly represent the true score function $s^{*}(x)=\nabla\log p_{t}(x)$. This error is defined as:
$\mathcal{E}^{\text{approx}}_{t}=\min_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t}}\left\|s_{\theta}(x)-\nabla\log p_{t}(x)\right\|^{2},$ (38)
where $\Theta$ denotes the parameter space of the neural network class.
To apply known approximation results for bounded target functions, we first analyze the probability that the score function $\nabla\log p_{t}(x)$ is bounded in norm. For a fixed threshold $\kappa>0$, consider
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\right).$ (39)
Using the law of total expectation, we have
$\mathbb{E}_{p}\left[\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\,\middle|\,x_{0}\right)\right].$ (40)
Now, conditioned on the data $p$, the forward process implies that $x_{t}$ follows a Gaussian distribution: $x_{t}\sim\mathcal{N}(\alpha_{t},\sigma^{2}_{t}{\cdot}I)$. Therefore, for fixed $x_{0}$
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\,\middle|\,x_{0}\right)=\mathbb{P}\left(\left\|\frac{x-\alpha_{t}}{\sigma_{t}}\right\|\geq\kappa\,\middle|\,x_{0}\right).$ (41)
The quantity $\frac{x-\alpha_{t}}{\sqrt{1-\alpha_{t}}}$ is Gaussian with mean 00 under this condition thus the expectation with respect to $p$ can be ignored. So we have
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\right)\leq Z_{\kappa,t}$ (42)
This implies
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\right)\leq Z_{\kappa}$ (43)
where the constant $Z_{\kappa,t}$ does not depend on $x_{{}_{0}}$. $Z_{\kappa}=\sup_{t\in[0,T]}Z_{\kappa,t}$ which ensures that this bound holds for all $t\in[0,T]$ denotes the tail probability of the standard normal distribution.
We can now fix $\kappa$ high enough such that that $Z_{\kappa}\leq\frac{\delta}{2}$, ensuring that with probability at least $1-\delta$, the score function is bounded in norm by $\kappa$. Now note that the expectation $E_{x\sim p_{t}}||s_{\theta}-\nabla{\log}p_{t}(x)||^{2}$ is bounded. Therefore, there exists an unbounded set $K$ such that $P(x\in K)\leq\frac{\delta}{2}$ and the complement of $K$ is a bounded set. Thus we get
$P(A\cup K)\leq P(A)+P(K)\leq{\delta}$ (44)
$P(\bar{A}\cap\bar{K})\geq 1-\delta$ (45)
Thus we have with probability at least $1-\frac{\delta}{2}$ that the function ${\nabla}{\log}p_{t}(x)$ is bounded and belongs to a bounded set. we invoke Theorem 2.5 from [36], which states that there exists a score function $s_{\tilde{\theta}}$ which is a neural network with ReLU activation such that.
$||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}\leq(W.D)^{-\frac{2}{d}}$ (46)
Now Theorem 6 of [41] states that for any neural network with ReLU activation $s_{\tilde{\theta}}$ there exists a neural network with a smooth activation such that.
$||s_{\tilde{\theta}}-s_{{\theta}}||_{\infty}\leq\epsilon$ (47)
for any $\epsilon>0$. Now setting $\epsilon=(W.D)^{-\frac{2}{d}}$ we get that there exists a neural network with a smooth activation function such that
$||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}$ $\leq||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}+||s_{\tilde{\theta}}-s_{{\theta}}||_{\infty}$ (48)
$\leq(W.D)^{-\frac{2}{d}}$ (49)
Since the bound holds for $||.||_{\infty}$ we can take expectation with respect to ${p_{t}}$ on both sides of Equation (49) to get the required result. ∎
D.2 Bounding the Statistical Error
Next, we bound the statistical error $\mathcal{E}_{\text{stat}}$. This error is due to estimating $s^{*}(x)$ from a finite number of samples rather than the full data distribution.
Recall Lemma 2 (Statistical Error.) Let $n$ denote the number of samples used to estimate the score function at each diffusion step. Then, for all $k\in[0,K]$, with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n}}}\right)$ (51)
Proof.
Let us define the population loss at time $t_{k}$ for $k\in[0,K]$ as
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{K}}}\left\|s_{\theta}(x,t)-\nabla\log p_{t}(x)\right\|^{2},$ (52)
where $s_{\theta}$ denotes the score function estimated by a neural network parameterized by $\theta$, and $x$ denotes samples at time $t$ used in Algorithm 1.
The corresponding empirical loss is defined as:
^$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t)-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (53)
Let $\theta^{a}_{k}$ and $\theta^{b}_{k}$ be the minimizers of $\mathcal{L}_{k}(\theta)$ and ^$\widehat{\mathcal{L}}_{t}(\theta)$, respectively, corresponding to score functions $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$. By the definitions of minimizers, we can write
$\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})$ $\leq\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})+\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})$ (54)
$\leq\underbrace{\left|\mathcal{L}_{k}(\theta^{b}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})\right|}_{\text{(I)}}+\underbrace{\left|\mathcal{L}_{k}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})\right|}_{\text{(II)}}.$ (55)
Note that right hand side of Equation (54) is less then the left hand side since we have added the quantity ^$\widehat{\mathcal{L}}_{t}(\theta^{a}_{t})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{t})$ which is strictly positive since $\theta^{b}_{k}$ is the minimizer of the function ^$\widehat{\mathcal{L}}_{t}(\theta)$ by definition. We then take the absolute value on both sides of Equation (55) to get
$|\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})|$ $\leq\underbrace{\left|\mathcal{L}_{k}(\theta^{b}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})\right|}_{\text{(I)}}+\underbrace{\left|\mathcal{L}_{k}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})\right|}_{\text{(II)}}.$ (56)
Note that this preserves the direction of the inequality since the left hand side of Equation (55) is strictly positive since $\theta^{a}_{k}$ is the minimizer of $\mathcal{L}_{k}(\theta)$ by definition.
We now bound terms (I) and (II) using standard generalization results. From Theorem 26.5 of [35], if the loss function is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta^{a}_{t},\theta^{b}_{t}\}$, then with probability at least $1-\delta$, we have
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\widehat{R}(\theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\forall~{}\theta\in\Theta^{\prime\prime}$ (57)
where $\widehat{R}(\theta)$ denotes the empirical Rademacher complexity of the function class restricted to $\Theta^{\prime\prime}$.
To ensure the boundedness of the loss, we define the event $A$ as
$A:=\left\{x:\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\right\}.$ (58)
As shown in Lemma 1, $\mathbb{P}(A)\geq 1-\delta$ for an appropriately large constant $\kappa$, based on the Gaussian nature of $x_{t}$. Let event $B$ denote the failure of the generalization bound
$B:=\left\{\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]>\widehat{R}(\Theta^{\prime\prime})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\right\}.$ (59)
From the above theorem, we know $\mathbb{P}(B)\leq\delta$ under the boundedness condition. Therefore, by the union bound, we have
$\mathbb{P}(A^{c}\cup B)$ $\leq\mathbb{P}(A^{c})+\mathbb{P}(B)\leq 2\delta,$ (60)
$\Rightarrow\mathbb{P}(A\cap B^{c})$ $\geq 1-2\delta.$ (61)
On this event $(A_{\kappa}\cap B^{c})$, both the score function is bounded and the generalization bound holds. Hence, with probability at least $1-2\delta$, we have:
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\widehat{R}(\theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}.$ (62)
Since $\Theta^{\prime\prime}$ is a finite class (just two functions). Note that Massart’s Lemma states that for a finite function class $\Theta^{{}^{\prime\prime}}$ we have
$\mathbb{E}_{\sigma}\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}\leq||f(\theta)||_{2}\leq d.||\Theta||$ (63)
We get the second inequality by denoting $d$ as the dimension of $f(\theta)$ and $||\Theta||$ be a quantity such that $||\Theta||=\sup_{\theta\in\Theta^{{}^{\prime\prime}}}||\theta||_{\infty}$. Also, $\sigma=\{{\sigma_{1}},{\sigma_{2}},{\cdots},\sigma_{n}\}$
where $\sigma_{i}$ are i.i.d random variables such that $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=\frac{1}{2}$.
Now, the definition of $\hat{R}(\theta)$ is $\hat{R}(\theta)=\frac{1}{m}\mathbb{E}_{\sigma}\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}$. Thus by applying Massart’s lemma to $\hat{R}(\theta)$ in Equation (62) we get with probability at least $1-2.\delta$
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\mathcal{O}\left(\frac{d.||\Theta||}{n}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}.$ (64)
Here $||\Theta||$ is a value such that $||\Theta||=\sup_{\theta\in\Theta^{{}^{\prime\prime}}}||\theta||_{\infty}$. Which implies that $\Theta$ is the norm of the maximum parameter value that can be attained in the finite function class $\Theta^{{}^{\prime\prime}}$. which yields that with probability at least $1-2.\delta$ we have
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}$ (65)
Finally, using the Polyak-Łojasiewicz (PL) condition for $\mathcal{L}_{k}(\theta)$, from Assumption 2, we have from the quadratic growth condition of PL functions the following,
$\|\theta^{a}_{t}-\theta^{b}_{t}\|^{2}\leq\mu\left|\mathcal{L}_{k}(\theta^{a}_{t})-\mathcal{L}_{k}(\theta^{b}_{t})\right|,$ (66)
and applying Lipschitz continuity of the score network with respect to parameters:
$\|s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\|^{2}$ $\leq L.\|\theta^{a}_{t}-\theta^{b}_{t}\|^{2}$ (67)
$\leq L.\mu\left|\mathcal{L}_{k}(\theta^{a}_{t})-\mathcal{L}_{k}(\theta^{b}_{t})\right|$ (68)
$\leq\mathcal{O}\left(d.\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (69)
Here $L$ is the lipschitz parameter of the neural network. Taking expectation with respect to $x\sim p_{t_{k}}$ on both sides completes the proof. For the final bound we replace $\delta$ with $\frac{\delta}{2}$ for notational consistency.
D.3 Bounding Optimization Error
The optimization error ($\mathcal{E}_{\text{opt}}$) accounts for the fact that gradient-based optimization does not necessarily find the optimal parameters due to limited steps, local minima, or suboptimal learning rates. This can be bounded as follows.
Recall Lemma 3 (Optimization Error.) Let $n$ be the number of samples used to estimate the score function at each diffusion step. If the step size satisfies $0\leq\eta\leq\frac{1}{\kappa}$, then under Assumptions 1, 2 and 3, for all $k\in[0,K]$, the optimization error due to imperfect minimization of the training loss satisfies with probability at least $1-\delta$:
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (70)
Proof.
Let $\mathcal{E}_{k}^{\text{opt}}$ denote the optimization error incurred when performing stochastic gradient descent (SGD), with the empirical loss defined by
^$\widehat{\mathcal{L}}^{i}_{k}(\theta)=\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (71)
The corresponding population loss is
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right],$ (72)
Thus, $\mathcal{E}_{t}^{\text{opt}}$ captures the error incurred during the stochastic optimization at each fixed time step $t$. We now derive upper bounds on this error.
From the smoothness of $\mathcal{L}_{k}(\theta)$ throgh Assumption 1, we have
$\mathcal{L}_{k}(\theta_{i+1})\leq\mathcal{L}_{k}(\theta_{i})+\langle\nabla\mathcal{L}_{k}(\theta_{i}),\theta_{i+1}-\theta_{i}\rangle+\frac{\kappa}{2}\|\theta_{i+1}-\theta_{i}\|^{2}.$ (73)
Taking conditional expectation given $\theta_{i}$, and using the unbiased-ness of the stochastic gradient ^$\nabla\widehat{\mathcal{L}}_{k}(\theta_{i})$, we get:
$\mathbb{E}[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}]\leq\mathcal{L}_{k}(\theta_{i})-\alpha_{t}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{\kappa\alpha^{2}_{t}}{2}\mathbb{E}[\|\nabla\widehat{\mathcal{L}}_{t}(\theta_{i})\|^{2}\mid\theta_{i}].$ (74)
Now using the variance bound on the stochastic gradients using Assumption 3, we have
$\mathbb{E}[\|\nabla\widehat{\mathcal{L}}_{k}(\theta_{i})\|^{2}\mid\theta_{i}]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$ (75)
Using this in the previous equation, we have that
$\mathbb{E}[\mathcal{L}_{k}(\theta_{t+1})\mid\theta_{i}]$ $\leq\mathcal{L}_{k}(\theta_{i})-\eta\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\frac{\kappa\eta^{2}}{2}\left(\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\sigma^{2}\right)$ (76)
$=\mathcal{L}(\theta_{i})-\left(\eta-\frac{\kappa\eta^{2}}{2}\right)\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\frac{\kappa\eta^{2}\sigma^{2}}{2}.$ (77)
Now applying the PL inequality (Assumption 2), $\|\nabla L(\theta_{i})\|^{2}\geq 2\mu\left(L(\theta_{i})-L^{*}\right)$, we substitute in the above inequality to get
$\mathbb{E}[\mathcal{L}(\theta_{t+1})\mid\theta_{i}]-\mathcal{L}^{*}$ $\leq\left(1-2\mu\left(\eta-\frac{\kappa\eta^{2}}{2}\right)\right)\left(\mathcal{L}(\theta_{i})-\mathcal{L}^{*}\right)+\frac{\kappa\eta^{2}\sigma^{2}}{2}.$ (78)
Define the contraction factor
$\rho\coloneqq 1-2\mu\left(\eta-\frac{\kappa\eta^{2}}{2}\right).$ (79)
Taking total expectation and defining $\delta_{t}=\mathbb{E}[L(\theta_{i})-L^{*}]$, we get the recursion:
$\delta_{t+1}\leq\rho\cdot\delta_{t}+\frac{\kappa\eta^{2}\sigma^{2}}{2}.$ (80)
When $\eta\leq\frac{1}{\kappa}$, we have
$\eta-\frac{\kappa\eta^{2}}{2}\geq\frac{\eta}{2}\Rightarrow\rho\leq 1-\mu\eta.$ (81)
Unrolling the recursion we have
$\delta_{t}\leq(1-\mu\eta)^{t}\delta_{0}+\frac{\kappa\eta^{2}\sigma^{2}}{2}\sum_{j=0}^{t-1}(1-\mu\eta)^{j}.$ (82)
Using the geometric series bound:
$\sum_{j=0}^{t-1}(1-\mu\eta)^{j}\leq\frac{1}{\mu\eta},$ (83)
we conclude that
$\delta_{t}\leq(1-\mu\eta)^{t}\delta_{0}+\frac{\kappa\eta\sigma^{2}}{2\mu}.$ (84)
Hence, we have the convergence result
$\mathbb{E}[\mathcal{L}(\theta_{i})-\mathcal{L}^{*}]\leq(1-\mu\eta)^{t}\delta_{0}+\frac{\kappa\eta\sigma^{2}}{2\mu}.$ (85)
From theorem 4.6 of [14] states that this result implies a sample complexity of $\mathcal{O}\left(\frac{1}{n}\right)$ where $n$ is the number of steps of the SGD algorithm performed. Note that the result from [14] requires bounded gradient and smoothness assumption which are satisfied by Assumption 1.
Note that $\hat{s}_{t_{k}}$ and $\hat{\theta}_{k}$ denote our estimate of the loss function and assosciated parameter obtained from the SGD. Also note that $\mathcal{L}^{*}$ is the loss function corresponding whose minimizer is the neural network $s_{t_{k}}^{a}$ and the neural parameter $\theta_{k}^{a}$ is our estimated score parameter. Thus applying the quadratic growth inequality.
$\|\hat{s}_{t_{k}}(x,t_{k})-s^{a}_{t_{k}}(x,t_{k})\|^{2}$ $\leq L.||\hat{\theta}_{k}-\theta^{k}_{a}||^{2}\leq||[\mathcal{L}(\theta_{k})-\mathcal{L}^{*}]||$ (86)
$\leq\mathcal{O}\left(\frac{1}{n}\right)$ (87)
From lemma 2 we have with probability $1-\delta$ that
$\|s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\|^{2}$ $\leq L.\|\theta^{a}_{t}-\theta^{b}_{t}\|^{2}$ (88)
$\leq L.\mu\left|\mathcal{L}_{k}(\theta^{a}_{t})-\mathcal{L}_{k}(\theta^{b}_{t})\right|$ (89)
$\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (90)
Thus we have with probability at least $1-\delta$
$||\hat{s}_{t}(x)-s^{b}_{t}(x)||^{2}$ $\leq 2||\hat{s}_{t_{k}}(x,t_{k})-s^{a}_{t_{k}}(x,t_{k})||+2.||s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})||$ (91)
$\leq\mathcal{O}\left({\log}\left(\frac{1}{n}\right)\right)+\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (92)
$\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (93)
Taking expectation with respect to $x\sim p_{t_{k}}$ on both sides completes the proof.