### v5

# Improved Sample Complexity For Diffusion Model Training Without Empirical Risk Minimizer Access

Abstract
Diffusion models have demonstrated remarkable performance in generating high-dimensional samples across domains such as vision, language, and the sciences. Although continuous-state diffusion models have been extensively studied both empirically and theoretically, discrete-state diffusion models, essential for applications involving text, sequences, and combinatorial structures, they remain significantly less understood from a theoretical standpoint. In particular, all existing analyses of discrete-state models assume access to an empirical risk minimizer. In this work, we present a principled theoretical framework analyzing diffusion models, providing a state-of-the-art sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$. Our structured decomposition of the score estimation error into statistical and optimization components offers critical insights into how diffusion models can be trained efficiently. This analysis addresses a fundamental gap in the literature and establishes the theoretical tractability and practical relevance of diffusion models.
1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks (Ulhaq & Akhtar, 2022; Bansal et al., 2023), text generation (Li et al., 2022), sequential data modeling (Tashiro et al., 2021), reinforcement learning and control (Zhu et al., 2023), and life sciences (Jing et al., 2022; Malusare & Aggarwal, 2024). For a more comprehensive exposition of applications, we refer readers to survey paper (Chen et al., 2024).
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. Some of the key works studying the sample complexity are summarized in Table 1. A key limitation of sample complexity analyses of diffusion models done thus far is the lack of the presence of finite-time sample complexity results under reasonable assumptions. This makes the theoretical analysis of diffusion models fall short of other machine learning areas such as reinforcement Learning (Kumar et al., 2023; Gaur et al., 2024), bi-level-optimization (Grazzi et al., 2023; Gaur et al., 2025) and graphical models (Fattahi et al., 2019; Tran et al., 2019). In this work we aim to bridge that gap and obtain a sample complexity results on the same footing as results from the aforementioned areas. The iteration complexity or convergence has been studied in Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b), while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function.
We note that works such as Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023); Chen et al. (2023) have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, Gupta et al. (2024) improved upon this by obtaining $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity without exponential dependence on data dimension. [^1] We note that Gupta et al. (2024) claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix A.. However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss, a significant restriction that was explicitly highlighted as an open problem in Gupta et al. (2024) itself. While this assumption is present in all prior works, it is an unrealistic assumption regardless.
In this paper, we do not make this assumptions and establish an improved state-of-the-art sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-4})$. This represents a key step toward bridging the gap between the theory and practice of diffusion models. Specifically, we address the following fundamental question.
How many samples are required for a sufficiently expressive neural network to estimate the score function well enough to generate high-quality samples using a DDPM algorithm?
Our analysis directly connects the quality of the learned score function to the total variation distance between the generated and target distributions, offering more interpretable and practically relevant guarantees. Additionally, our work accounts for the unavailability of the empirical risk minimizer. Using our novel analysis of the score estimation error, we obtain the sample complexity bounds without exponential dependence on the data-dimension. Our principled analysis accounts for the statistical and optimization errors while not assuming access to the empirical risk minimizer of the score estimation loss, and achieves state-of-the-art sample complexity bounds.
Reference
Sample
Complexity
Empirical Risk
Minimizer Assumption
Zhang et al. (2024) $\tilde{O}\left(\epsilon^{-d}\right)$ Yes
Wibisono et al. (2024) $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes
Oko et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Chen et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Gupta et al. (2024)1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes
This work $\tilde{O}(\epsilon^{-4})$ No
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. For further details on how the sample complexity bounds are derived for Gupta et al. (2024), see Appendix A.
The statistical error occurs due to the finite sample size used to obtain the score estimate. Existing methods used to bound statistical errors assume bounded loss functions, which is not true in the case of diffusion models. We thus use a novel analysis that uses the conditional normality of the score function to obtain upper bounds for the statistical error.
Finally, the optimization error occurs due to a finite number of SGD steps during the estimation of the score function. It is precisely the error that was not accounted for in the previous works due to the assumption that they have access to the empirical risk minimizer. We use the quadratic growth property implied by the Polyak-Łojasiewicz (PL) condition assumed in Assumption 2 and a novel recursive analysis of the error at each stochastic gradient descent (SGD) step to upper bound this error.
The main contributions of our work are summarized as:
  • Finite time sample complexity bounds. We derive state-of-the-art sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$ for score-based diffusion models, without exponential dependence on the data dimension or neural network parameters. Our analysis avoids the unrealistic assumptions used in prior works such as access to an empirical loss minimizer.
  • Principled error decomposition. We propose a structured decomposition of the score estimation error into approximation, statistical, and optimization components, enabling the characterization of how each factor contributes to sample complexity.
Unconditional and Conditional Diffusion Models. Diffusion models have emerged as leading frameworks across vision, audio, and scientific domains. Foundational works such as Sohl-Dickstein et al. (2015) and Ho et al. (2020) introduced and refined Denoising Diffusion Probabilistic Models (DDPMs), enabling high-quality sample generation. Subsequent advances include improved noise schedules (Nichol & Dhariwal, 2021), score-based SDE formulations (Song et al., 2021), and efficient latent-space generation via Latent Diffusion Models (LDMs) (Rombach et al., 2022). Conditional diffusion models extend these techniques for guided generation tasks, with applications in time-series Tashiro et al. (2021), speech Huang et al. (2022), and medical imaging Dorjsembe et al. (2023). Conditioning mechanisms range from classifier-based Dhariwal & Nichol (2021) to classifier-free guidance Ho & Salimans (2022), which enabled text-to-image models like Imagen Saharia et al. (2022) and Stable Diffusion Rombach et al. (2022). Recent innovations focus on adaptive control Castillo et al. (2025), compositionality Liu et al. (2023), and multi-modal conditioning Avrahami et al. (2022).
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in Chen et al. (2022) showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b). In particular, Benton et al. (2024); Li et al. (2024b) establishes iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates Li et al. (2024a); Liang et al. (2025b); Dou et al. (2024). Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in Li & Yan (2024); Huang et al. (2024); Liang et al. (2025a). In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the errors introduced by the neural network approximation, data sampling, and optimization are also accounted for.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in Chen et al. (2023); Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023). However, these bounds are exponential in the data dimension. Further, authors of Gupta et al. (2024) use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
2 Preliminaries and Problem Formulation
We begin by outlining the theoretical basis of score-based diffusion models. In particular, we adopt the continuous-time stochastic differential equation (SDE) framework, which provides a principled basis for modeling the generative process. We then outline its practical discretization and formally define our problem.
Score-based generative models enable sampling from a complex distribution $p_{0}$ by learning to reverse a noise-adding diffusion process. This approach introduces a continuous-time stochastic process that incrementally perturbs the data distribution into a tractable distribution (typically Gaussian), and then seeks to reverse that transformation.
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process (Øksendal, 2003), defined by the following SDE
$dx_{t}=-x_{t}dt+\sqrt{2}dB_{t},\quad x_{0}\sim p_{0},x\subset\mathbb{R}^{d},$ (1)
where $B_{t}$ denotes standard Brownian motion. The solution of this SDE in closed form is given by
$x_{t}\sim e^{-t}x_{0}+\sqrt{1-e^{-2t}}{\epsilon},\quad\text{with }\epsilon\sim\mathcal{N}(0,I).$ (2)
As $t\to\infty$, the process converges to the stationary distribution $\mathcal{N}(0,I)$. Let $p_{t}$ denote the marginal distribution of $x_{t}$. This defines a continuous-time smoothing of the data distribution, where $p_{t}$ becomes increasingly Gaussian over time.
This is typically achieved using stochastic time-reversal theory (Anderson, 1982), which yields a corresponding reverse-time SDE as follows.
$dx_{T-t}=\left(x_{T-t}+2\nabla\log p_{T-t}(x_{T-t})\right)dt+\sqrt{2}\,dB_{t},$ (3)
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions. In practice, the backward process is run up to a fixed time point $t_{0}$ known as the early stopping time and not $t=0$. This is done in order to improve performance and training speed (Lyu et al., 2022; Favero et al., 2025).
The continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework (Ho et al., 2020), where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix B. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
Problem formulation: Let the score function be approximated using a parameterized family of neural networks $\mathcal{F}_{\Theta}=\{s_{\theta}:\theta\in\Theta\}$, where each $s_{\theta}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ is represented by a neural network of depth $D$ and width $W$ with smooth activation functions. Given $n_{k}$ i.i.d. samples $\{x_{i}\}_{i=1}^{n}$ from the data distribution $p_{t_{k}}$, the score network is trained by minimizing the following time-indexed loss:
$\mathcal{L}_{k}(\theta):=\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}\right].$ (4)
Objective. Our goal is to quantify how well the learned generative model $\hat{p}_{t_{0}}$ approximates the true data distribution $p$ in terms of total variation (TV) distance. Specifically, we aim to show the number of samples needed so that with high probability, the TV distance $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ is bounded by $\mathcal{O}(\epsilon)$, where $\epsilon$ is the $L^{2}$ estimation error of the score function. This reduces the generative performance analysis to establishing tight sample complexity bounds on the score estimation error, using only $n$ i.i.d. samples from $p$. We additionally define the following probability distributions:
$p_{t_{0}}:=$ Distribution obtained after backward process till time $t_{0}$ steps starting form $p_{T}$
$p^{dis}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
at discretized time steps
$\tilde{p}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
at discretized time steps using the estimated score functions
$\hat{p}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $\mathcal{N}(0,I)$
at discretized time steps using the estimated score functions
where $t_{0}$ denotes the early stopping time.
3 Sample Complexity of Diffusion Models
In this section, we derive explicit sample complexity bounds for diffusion-based generative models. By leveraging tools from stochastic optimization and statistical learning theory, we provide bounds on the number of data samples required to accurately estimate the time-dependent score function $s_{t}(x):=\nabla\log p_{t}(x)$ across the forward diffusion process. Note that accurate score estimation is critical for ensuring high-quality generation while sampling through the reverse-time SDE.
We first state the assumptions required throughout this work.
Assumption 1 (Bounded Second Moment Data Distribution.).
[…]
Some works that analyze the convergence of score-based diffusion models, such as Chen et al. (2022); Oko et al. (2023), assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to be sub-Gaussian, making our results applicable to a significantly broader class of distributions.
Assumption 2 (Polyak - Łojasiewicz (PL) condition.).
The loss $\mathcal{L}_{k}(\theta)$ for all $k\in[0,K]$ satisfies the Polyak–Łojasiewicz condition, i.e., there exists a constant $\mu_{t}>0$ such that
$\frac{1}{2}\|\nabla\mathcal{L}_{k}(\theta)\|^{2}\geq\mu_{t}\left(\mathcal{L}_{k}(\theta)-\mathcal{L}_{k}(\theta^{*})\right),\quad\forall~\theta\in\Theta,$ (5)
where $\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}_{k}(\theta)$ denotes the global minimizer of the population loss.
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses (Liu et al., 2022). Prior works such as Gupta et al. (2024) and Block et al. (2020) implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in Gupta et al. (2024) and the definition of $\hat{f}$ in Theorem 13 of Block et al. (2020)). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.
In contrast, the PL condition allows us to derive sample complexity bounds under realistic optimization dynamics, without requiring exact ERM solutions. To our knowledge, this is the first theoretical analysis of score-based generative models that explicitly accounts for inexact optimization, addressing a key gap in existing literature. Additionally, we establish convergence guarantees with both constant and decreasing step sizes.
Assumption 3 (Approximation error of the Class of Neural Networks).
For all $t\in[0,T]$, there exists a neural network parameter $\theta\in\Theta$ such that
[…]
This error is independent of the sampling algorithm, and describes the error due to neural network parametrization. In learning theory, it is common to treat the approximation error of a model class as a constant so that analyzes can focus on the estimation/ optimization terms dependent on the sample. This convention appears in standard excess-risk decompositions for fixed hypothesis classes (Shalev-Shwartz & Ben-David, 2014). In PAC-Bayesian analyses, approximation errors are denoted by a constant once the class is fixed (Mai, 2025). In (NTK/RKHS) analyses of neural networks, where it is assumed the target function lies in, or is well approximated by the specified function class, the misspecification error is represented as a constant term (Bing et al., 2025). In reinforcement learning algorithm analysis such as policy gradient, a task-dependent “inherent Bellman” or function-approximation error that remains constant while deriving performance rates (Mondal & Aggarwal, 2024; Fu et al., 2021; Gaur et al., 2024; Ganesh et al., 2025). Note that Gupta et al. (2024) also makes the same assumption implicitly, but assumes this constant to be zero (in Assumption A2).
Note that in certain works such as (Jiao et al., 2023), it is shown that the network size has to exponential in data dimension in order to achieve a small approximation error. However, in practice that would require an impractically large neural network size. In practice neural network size is of the same order as the data dimension. Thus for a fixed neural network size that we assume in this work, it makes sense to assume the approximation error as a constant.
Assumption 4 (Smoothness and bounded gradient variance of the score loss.).
For all $k\in[0,K]$, the population loss $\mathcal{L}_{k}(\theta)$ is $\kappa$-smooth with respect to the parameters $\theta$, i.e., for all $\theta,\theta^{\prime}\in\Theta$
$\|\nabla\mathcal{L}_{k}(\theta)-\nabla\mathcal{L}_{k}(\theta^{\prime})\|\leq\kappa\|\theta-\theta^{\prime}\|.$ (7)
We assume that the estimators of the gradients ${\nabla}\mathcal{L}_{k}(\theta)$ have bounded variance.
[…]
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function (Liu et al., 2022). Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models (Salimans & Ho, 2022; Liu et al., 2022). Note that in no prior works were such assumptions stated since they assumed access to the empirical risk minimizer.
Theorem 1 (Total Variation Distance Bound).
Let $p_{t_{0}}$ denote the distribution obtained by the backward process till time $t_{0}$ starting form $p_{T}$, and $\hat{p}_{t_{k}}(x)$ be the distribution generated by the backward process at discretized time steps $\{t_{k}\}$ using the estimated score functions $\hat{s}_{T-t_{k}}(x)$ where $k\in[0,K]$. Let $d$ be the data dimension, and $n_{k}$ be the number of samples for score estimation at time step $t_{k}$.
Assume that the data distribution satisfies Assumption 1, the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 3,4 for all $k\in[0,K]$ and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$. Further assume
$n_{k}={\Omega}\left(W^{2D}{\cdot}d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (9)
Then, with probability at least $1-\delta$, the total variation distance between the $p_{t_{0}}$ and $\hat{p}_{t_{0}}$ satisfies
$TV(p_{t_{0}},\hat{p}_{t_{0}})\leq\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon)+\epsilon_{approx}$ (10)
Furthermore, by setting $T=\Omega\left(\log\left(\frac{1}{\epsilon}\right)\right)$ and $K=\Omega(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (11)
with probability at least $1-\delta$.
Theorem 1 establishes that the total variation distance between the true data distribution and the diffusion model’s output can be made arbitrarily small specifically, $\tilde{\mathcal{O}}(\epsilon)$ by properly scaling model capacity and algorithmic parameters. To the best of our knowledge, these are the only known sample complexity bounds for score-based diffusion models, improving upon the prior results as discussed in the introduction without assuming access to empirical risk minimizer for the score estimation loss. Our results highlight an inherent trade-off between the sample complexity bounds and the assumption of access to the empirical risk minimizer.
Usage of $p_{t_{0}}$ instead of $p_{0}$ in Theorem 1: We have shown that the estimated distribution $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in total variation (TV) to $p_{t_{0}}$, where $p_{t_{0}}$ denotes the data distribution pushed forward by $t_{0}$ steps of the forward process. We do not claim that $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in TV to the true data distribution $p_{0}$ (i.e., we do not bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$), because doing so would require additional assumptions on $p$. For example, Fu et al. (2024) (in Lemma D.5) assumes a sub-Gaussian data distribution to show that $\mathrm{TV}(p_{0},p_{t_{0}})\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)$. We also note that all other works listed in Table 1 similarly provide upper bounds on $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$, not on $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$.
However, it is to be noted that using the sub-Gaussian assumption, our analyis can be extended to a bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$ via the triangle inequality:
$\mathrm{TV}(p_{0},\hat{p}_{t_{0}})\;\leq\;\mathrm{TV}(p_{0},p_{t_{0}})\;+\;\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$
We formally present the data assumption and the resulting theorem as follows
Assumption 5 (Sub-Gaussian Data Distribution.).
The data distribution $p_{0}$ of the data variable $x_{0}$ has an absolutely continuous CDF, is supported on a continuous set $\Gamma\in\mathbb{R}^{d}$, and there exists a constant $0<C_{2}<\infty$ such that for every $t\geq 0$ we have $P(|x_{0}|\geq t)\leq 2\cdot{\exp}^{-\frac{t^{2}}{C^{2}_{2}}}$.
Theorem 2 (Total Variation Distance Bound Under Sub-Gaussian Assumption).
Assume that the data distribution satisfies Assumption 5, the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 3,4 for all $k\in[0,K]$ and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$. Further assume
$n_{k}={\Omega}\left(W^{2D}{\cdot}d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (12)
Then, with probability at least $1-\delta$, the total variation distance between the $p_{0}$ and $\hat{p}_{t_{0}}$ satisfies
$TV(p_{0},\hat{p}_{t_{0}})$ $\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon)+\epsilon_{approx}$ (13)
Furthermore, by setting $t_{0}=\mathcal{O}(\epsilon^{-2})$, $T=\Omega\left(\log\left(\frac{1}{\epsilon}\right)\right)$ and $K=\Omega(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p_{0},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (14)
with probability at least $1-\delta$.
Proof of Theorem 1.
Recall that $\hat{p}_{t_{0}}$ is derived via score-based sampling, so using the triangle inequality repeatedly to decompose the TV distance between the true distribution $p_{t_{0}}$ and $\hat{p}_{t_{0}}$ as follows
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},\hat{p}_{t_{0}},)$ (15)
The bounds on $\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})$ and $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ follow from Lemma B.4 of Gupta et al. (2024) and Proposition 4 of Benton et al. (2024), respectively to get
$\mathrm{TV}((p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (16)
Note that we have used results from Gupta et al. (2024) and Benton et al. (2024) which assume a bounded second moment for the data distribution. This is satisfied by Assumption 1. Now from lemma 4, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ is upper bounded as follows
$\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})\leq\frac{1}{2}\sqrt{\sum_{k=0}^{K}E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}(t_{k+1}-t_{k})}$ (17)
In order to upper bound, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ we denote $A(k)$ as
$A(k):=E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}dx$ (18)
Therefore, bounding the TV distance between $p$ and $\hat{p}_{t_{0}}$ translates to bounding the cumulative error in estimating the score function at different time steps. We now focus on bounding this term. Specifically, we have that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\frac{1}{2}\sqrt{\sum_{k=0}^{K}A_{k}{\cdot}(t_{k+1}-t_{k})dx}+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (19)
Now, for each time step $k$, we decompose the total score estimation error, denoted by $A(k)$, into three primary components: approximation error, statistical error, and optimization error. Each of these error corresponds to a distinct aspect of learning the reverse-time score function in a diffusion model as described below.
$\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right]$ $\leq 4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|s^{a}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{approx}}_{k}}$
$\quad+4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{stat}}_{k}}$
$\quad+4\underbrace{\mathbb{E}_{x\sim p_{t_{k}}(x)}\left[\left\|\hat{s}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\right\|^{2}\right]}_{\mathcal{E}^{\mathrm{opt}}_{k}},$ (20)
where, we define the parameters
$\theta^{a}_{k}$ $=\arg\min_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t}(x,t_{k})\|^{2}\right],$ (21)
$\theta^{b}_{k}$ $=\arg\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i},t_{k})\right\|^{2}$ (22)
and denote $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$ as the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$ respectively. Approximation error $\mathcal{E}^{\mathrm{approx}}_{k}$ captures the error due to the limited expressiveness of the function class $\{s_{\theta}\}_{\theta\in\Theta}$. The statistical error $\mathcal{E}^{\mathrm{stat}}_{k}$ is the error from using a finite sample size. Finally, the optimization error $\mathcal{E}^{\mathrm{opt}}_{k}$ is due to not reaching the global minimum during training.
One of our key contributions lies in rigorously bounding each of these error components and showing how their interplay governs the overall generative error. In particular, we derive novel bounds that explicitly capture the dependencies on sample size, neural network capacity, and optimization parameters, without any assumption on the access to the empirical risk minimizer of the score estimation loss. We formalize these results in the following lemmas. Detailed proofs are deferred to Appendices C.1, and C.2, respectively.
Lemma 1 (Approximation Error).
$\mathcal{E}_{k}^{\mathrm{approx}}$ is defined as follows
$\mathcal{E}_{k}^{\mathrm{approx}}=\text{min}_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t}(x,t_{k})\|^{2}\right]$ (23)
Then, under Assumption 3 for all $k\in[{0},K]$, we have
$\mathcal{E}^{\mathrm{approx}}_{k}\leq\epsilon_{approx}$ (24)
This result directly follows from Assumption 3 and the definition of $\mathcal{E}^{\mathrm{approx}}_{k}$.
Lemma 2 (Statistical Error).
Let $n_{k}$ denote the number of samples used to estimate the score function at each diffusion step $t_{k}$. If the data distribution satisfies the Assumption 1 and the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 for all $k\in[{0},K]$, then with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq{\mathcal{O}}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)$ (25)
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Existing analyses of statistical errors, such as those given in Shalev-Shwartz & Ben-David (2014), only work when the loss function is bounded. This is not the case for diffusion models. Thus, we use a novel analysis that uses the conditional normality of the score function as well as the bounded second moment property of the data variable in Assumption 1 to obtain the upper bound on the statistical error. The details of the analysis are given in Appendix C.1.
Lemma 3 (Optimization Error).
Let $n_{k}$ be the number of samples used to estimate the score function at diffusion step $t_{k}$}. Assume that the score loss function $\mathcal{L}_{k}(\theta)$ satisfies the Assumptions 2 and 4, for all $k\in[0,K]$, and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$, then with probability at least $1-\delta$
$\mathcal{E}^{\mathrm{opt}}_{k}\leq{\mathcal{O}}\left(W^{D}.d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right).$ (26)
We leverage assumptions 2, 4 , alongside our unique recursive at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function. This is the first analysis of diffusion models to explicitly account for this error. All prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly. The details of the analysis are given in Appendix C.2.
Combining the decomposition in equation 20 along with Lemmas 13, we obtain the following bound on $A(k)$ ( equation 18) with probability at least $1-\delta$
$A(k)$ $\leq\mathcal{O}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left(W^{D}.d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right)+\epsilon_{approx}$ (27)
$\leq\mathcal{O}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\epsilon_{approx},$ (28)
where in the second inequality we combine the first two terms appropriately. Setting the sample size
$n_{k}={\Omega}\left(W^{2D}.d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (29)
we ensure that $A(k)\leq\frac{\epsilon^{2}}{\sigma_{k}^{2}}=\frac{\epsilon^{2}}{1-e^{-2(T-t_{k})}}$ for all $k\in\{0,\dots,K-1\}$. Summing over all time steps, we obtain with probability at least $1-\delta$
$\sum_{k=0}^{K}A(k)(t_{k+1}-t_{k})$ $\leq\sum_{k=0}^{K}\frac{\epsilon^{2}}{1-e^{-2(T-t_{k})}}\leq\int_{0}^{T}\frac{\epsilon^{2}}{1-e^{-2(T-t)}}dt\leq\epsilon^{2}T.$ (30)
Note that the term $\left(\log^{2}\left(\frac{4K}{\delta}\right)\right)$ appears in the upper bound for $n_{k}$ in equation 29 since we have to take a union bound for Lemma 2 and Lemma 3 and then take a union bound over $K$ discretization steps. Substituting this bound into equation 18, and then substituting the result into equation 16, we obtain that with probability at least $1-\delta$.
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon\cdot\sqrt{T})+\epsilon_{approx}$ (31)
Finally, by choosing $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we conclude that with probability at least $1-\delta$
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (32)
completing the proof of Theorem 1. ∎
In summary, our work provides a principled decomposition of the errors in score-based generative models, highlighting how each component contributes to the overall sample complexity. This leads to the first finite sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$ for diffusion models without assuming access to the empirical minimizer of the score estimation function.
4 Conclusion and Future Work
In this work, we investigate the sample complexity of training diffusion models via score estimation using neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$, which, to our knowledge, is the first such result that does not assume access to an empirical risk minimizer of the score estimation loss. Notably, our bound does not depend exponentially on the number of neural network parameters. For comparison, the best-known existing result achieves a bound of $\widetilde{\mathcal{O}}(\epsilon^{-5})$, but it crucially assumes access to an ERM. All prior results establishing sample complexity bounds for diffusion models have made this assumption. Our contribution is the first to establish a sample complexity bound for diffusion models under the more realistic setting where exact access to empirical risk minimizer of the score estimation loss is not available.
While our analysis focuses on unconditional distributions, extending these guarantees to conditional settings remains an important direction for future work.
Appendix A Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in Gupta et al. (2024), and show how combining their results with those of Block et al. (2020) leads to an alternative bound.
A.1 Sample Complexity of Gupta et al. (2024)
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in Gupta et al. (2024). A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of Gupta et al. (2024) is as follows. We re-iterate this Theorem here.
Theorem C.2 Gupta et al. (2024). Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
$\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|s_{\theta}(x)-s_{t_{k}}(x)\right\|_{2}^{2}\right]\leq\frac{\delta\cdot\epsilon^{3}}{CK^{2}\sigma_{T-t_{k}}^{2}}\cdot\frac{1}{\log\frac{d+m_{2}+1/m_{2}}{\gamma}}$ (33)
for sufficiently large constant $C$, then consider the score functions trained from
$m>\widetilde{\mathcal{O}}\left(\frac{K(d+\log\frac{1}{\delta})\cdot PD}{\epsilon^{3}}\cdot\log\left(\frac{\max(m_{2},1)\cdot\Theta}{\delta}\right)\cdot\log\left(\frac{m_{2}+1/m_{2}}{\gamma}\right)\right).$ (34)
i.i.d. samples of $q$, with $1-\delta$ probability, DDPM can sample from a distribution $\epsilon$-close in TV to a distribution $\gamma m_{2}$-close in 2-Wasserstein to $q$ in $N$ steps.
Note that Lemma B.6 of Gupta et al. (2024) states that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\delta}+\mathcal{O}\left(\frac{1}{\sqrt{N}}\right)+{\epsilon}\cdot{\sqrt{T}}+\mathcal{O}(\exp^{-T})$ (35)
Here $p$ and $\hat{p}$ are the true and learned data distributions, respectively.
In order to achieve $\mathrm{TV}(p,\hat{p})\leq\epsilon$ we have to set $\delta=\epsilon$. This would imply $N=\mathcal{O}(\epsilon^{-2})$ and $T=\mathcal{O}\left(\frac{1}{\epsilon}\right)$. Putiing this value of $N$ in Equation equation 34 we obtain that for
$m>\widetilde{O}\left(\frac{(d+\log(1/\delta))\cdot PD}{\epsilon^{5}}\cdot\log\left(\frac{\Theta}{\epsilon}\right)\right).$ (36)
we have with probability at least $1-\epsilon$
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\epsilon$ (37)
This reveals a discrepancy between the reported sample complexity and the actual bound derived above, highlighting that the true complexity is significantly higher than what was originally reported.
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike Gupta et al. (2024), our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
Appendix B Score Estimation Algorithm
In this section, we provide a detailed description of the algorithm used for estimating the score function in diffusion models.
Algorithm 1 Denoising Diffusion Probabilistic Model (DDPM)
1:Input: Dataset $\mathcal{D}$, timesteps $T$, stop time $t_{\text{stop}}$, schedule $\{\beta_{t}\}_{t=1}^{T}$, network $\epsilon_{\theta}$, learning rate $\eta$, iterations $K$
2: Precompute: $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ Training (Score Estimation)
3:for $k=1$ to $K$ do
4:  for $i=1$ to $n$ do
5:   Sample $x_{j}\sim\mathcal{D}$, $\epsilon_{k}\sim\mathcal{N}(0,I)$ for $i=1,\dots,n$
6:   $x_{t_{i}}=e^{-{t_{k}}}x_{i}+\sqrt{1-e^{-2t_{k}}}\epsilon_{i}$
7:   Compute loss: $\hat{L}(\theta)=\|\epsilon_{i}-\epsilon_{\theta}(x_{t_{i}},t_{i})\|^{2}$
8:   Update $\theta\leftarrow\theta-\eta_{i}\cdot\nabla_{\theta}\hat{L}(\theta)$
9:  end for
10:end forSampling
11: Sample $x_{T}\sim\mathcal{N}(0,I)$
12:for $t=T$ down to $t_{\text{stop}}+1$ do
13:  $z\sim\mathcal{N}(0,I)$ if $t>1$ else $z=0$
14:  $\hat{\epsilon}=\epsilon_{\theta}(x_{t},t)$
[…]
16:  $x_{t-1}=\tilde{\mu}_{t}+\sqrt{\beta_{t}}\cdot z$
17:end for
18:Return $x_{t_{\text{stop}}}$
Appendix C Proofs of Intermediate Lemmas
In this section, we present the proofs of intermediate lemmas used to bound the statistical error and optimization error in our analysis.
C.1 Bounding the Statistical Error
Proof.
Let us define the population loss at time $t_{k}$ for $k\in[0,K]$ as
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{k}}}\left\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2},$ (38)
where $s_{\theta}$ denotes the score function estimated by a neural network parameterized by $\theta$, and $x$ denotes samples at time $t$ used in Algorithm 1. The corresponding empirical loss is defined as:
$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t_{k}}(x_{i})\right\|^{2}.$ (39)
Let $\theta^{a}_{k}$ and $\theta^{b}_{k}$ be the minimizers of $\mathcal{L}_{k}(\theta)$ and $\widehat{\mathcal{L}}_{t}(\theta)$, respectively, corresponding to score functions $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$. By the definitions of minimizers, we can write
$\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})$ $\leq\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})+\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})$ (40)
$\leq\underbrace{\left|\mathcal{L}_{k}(\theta^{b}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})\right|}_{\text{(I)}}+\underbrace{\left|\mathcal{L}_{k}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})\right|}_{\text{(II)}}.$ (41)
We now bound terms (I) and (II) using generalization results. From Lemma 5 (Theorem 26.5 of Shalev-Shwartz & Ben-David (2014)), if the loss function $\widehat{\mathcal{L}}(\theta)$ is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta_{k}^{a},\theta_{k}^{b}\}$, then with probability at least $1-\delta$, we have
$\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\leq\widehat{R}(\Theta^{{}^{\prime\prime}})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\forall~\theta\in\Theta^{\prime\prime}$ (42)
where $\widehat{R}(\Theta^{{}^{\prime\prime}})$ denotes the empirical Rademacher complexity of the function class restricted to $\Theta^{\prime\prime}$. Now since $x$ is not bounded, this result does not hold. We then define the following two functions
$\mathcal{L^{\prime}}_{k}(\theta)=\mathbb{E}_{x\sim\mu_{t_{k}}}\left\|v_{\theta}(x,t_{k})-v_{t_{k}}(x)\right\|^{2},$ (43)
and
$\widehat{\mathcal{L^{\prime}}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|v_{\theta}(x_{i},t_{k})-v_{t_{k}}(x_{i})\right\|^{2}.$ (44)
where we define the functions
[…]
and
$\left(v_{\theta}(x,t)\right)_{j}=\begin{cases}\left(s_{\theta}(x,t)\right)_{j}&\text{if }|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}\leq\kappa\\ 0&\text{if }|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}\geq\kappa\end{cases}$ (46)
Here $\left(v_{t}(x)\right)_{j}$$\left(\nabla\log p_{t}(x)\right)_{j}$, $\left(v_{\theta}(x,t)\right)_{j}$ and $\left(s_{\theta}(x,t)\right)_{j}$ denote the $j^{th}$ co-ordinate of $v_{t}(x)$, $\left(\nabla\log p_{t}(x)\right)$, $v_{\theta}(x,t)$ and $s_{\theta}(x,t)$ respectively. Further, $|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}$ denotes the $j^{th}$ co-ordinate of the $i^{th}$ sample of the score function in $\hat{\mathcal{L}_{k}}(\theta)$ which is given by ${\log}p_{t}(x)=|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|$.
Note that the functions $v_{t}(x)$ and $v_{\theta}(x,t)$ are uniformly bounded. Thus using Theorem 26.5 of Shalev-Shwartz & Ben-David (2014) we have with probability at least $1-\delta$,
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\widehat{R}(\Theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (47)
Since $\Theta^{\prime\prime}=\{\theta_{a},\theta_{b}\}$ is a finite class (just two functions). We can apply Lemma E.5 to bound the empirical Rademacher complexity $\widehat{R}(\theta)$ in terms of the Rademacher complexity $R(\theta)$ of the function class $\Theta^{\prime\prime}$. Since $\widehat{R}(\theta)=\frac{1}{m}\mathbb{E}_{\sigma}\left[\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}\right]$, applying Lemma E.5, we have with probability at least $1-2\delta$
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\mathcal{O}\left(\frac{d\cdot W^{D}}{n}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (48)
This yields that with probability at least $1-\delta$ we have
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}$ (49)
From this we have
$\left[\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (50)
Now consider the probability of the event
$A_{i,j}=\left\{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right\}$ (51)
Where $\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}$ denotes the $k^{th}$ co-ordinate of of the $i^{th}$ sample of the score function given by $\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|$ We have the probability of this event upper bounded as
$P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)$ $=\mathbb{E}_{z}P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)$ (52)
$\leq\exp\left(-{\kappa^{2}(1-e^{-t})}\right)$ (53)
$\leq\exp\left(-{\kappa^{2}}\right)$ (54)
Setting $\kappa={{\log}\left(\frac{dn}{\delta}\right)}$, we have
$P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)\leq\frac{\delta}{dn}$ (55)
If we denote the event $A=\{\widehat{L}^{\prime}(\theta)=\widehat{L}(\theta)\}$, then by union bound we have $P(A)=P(\cup_{i,j}A_{i,j})\leq\sum_{i,j}P(A_{i,j})\leq\delta$. Let event $B$ denote the failure of the generalization bound, i.e.,
$B:=\left\{\left|\mathcal{L^{\prime}}_{t}(\theta)-\widehat{\mathcal{L^{\prime}}}_{t}(\theta)\right|>\widehat{R}(\Theta^{\prime\prime})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\right\}.$ (56)
From above, we know $\mathbb{P}(B)\leq\delta$ under the boundedness condition. Therefore, by the union bound, we have
$\mathbb{P}(A\cup B)$ $\leq\mathbb{P}(A)+\mathbb{P}(B)\leq 2\delta,$ (57)
$\implies\mathbb{P}(A^{c}\cap B^{c})$ $=1-P(A\cup B)\geq 1-2\delta.$ (58)
On this event $(A^{c}\cap B^{c})$, we have $\widehat{\mathcal{L}}^{\prime}(\theta)=\widehat{\mathcal{L}}(\theta)$. Hence, with probability at least $1-2\delta$, we have
$|\mathcal{L}_{k}(\theta_{k}^{b})-\mathcal{L}_{k}(\theta_{k}^{a})|$ $\leq\left|\mathcal{L}_{k}(\theta_{k}^{b})-\widehat{\mathcal{L}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L}}_{t}(\theta_{k}^{a})\right|.$ (59)
$\leq\left|\mathcal{L}_{k}(\theta_{k}^{b})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$ (60)
$=\left|\mathcal{L}_{k}(\theta_{k}^{b})-{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$
$\quad+\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{b})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$ (61)
$\leq\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{a})-{\mathcal{L}}_{t}(\theta_{k}^{a})\right|+\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{b})-{\mathcal{L}}_{t}(\theta_{k}^{b})\right|+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (62)
In order to bound $|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|$ we have the following
$|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|$ $\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}|(s_{\theta}(x,t_{k}))-({\nabla}{\log}p_{t_{k}}(x))|_{j}^{2}-\mathbb{E}_{x\sim u_{t}}|(v_{t}(x))_{k}-(v_{\theta}(x,t))|_{j}^{2}$ (63)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|({\nabla}{\log}p_{t_{k}})-(s_{\theta}(x,t_{k}))|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (64)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}-(s_{\theta}(x,t))_{k}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (65)
$\leq 2\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$
$+\sum_{j=1}^{d}2\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left((s_{\theta}(x,t))_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (66)
$\leq 2\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$
$+\sum_{j=1}^{d}C_{\Phi^{{}^{\prime\prime}}}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\textbf{1}_{\Big|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Big|\geq\kappa}\right)$ (67)
$\leq\left(\frac{4+2C_{\Phi^{{}^{\prime\prime}}}}{{\sigma_{t}}^{2}}\right)\underbrace{\sum_{k=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)}_{I}$
$+\frac{2}{{\sigma_{t}}^{2}}\underbrace{\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{0}|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)}_{II}$ (68)
We get Equation equation 66 from Equation equation 65 by using the identity $(a-b)^{2}\leq 2|a|^{2}+2|b|^{2}$. We get Equation equation 67 from Equation equation 66 by using Lemma 8. We get Equation equation 68 from Equation equation 67 by using the identity $(a-b)^{2}\leq 2|a|^{2}+2|b|^{2}$ again. Now we separately obtain upper bounds for the terms $I$ and $II$ as follows.
$\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (69)
$=\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (70)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}|x_{0}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (71)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{j}\sim(u_{t})_{k}|x_{0}}\left(|x_{j}|^{2}\Bigg|\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)P\left(\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|\geq\kappa\Bigg|x_{0}\right)$ (72)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{0}\sim(u_{t})_{j}|x_{0}}\left(|x_{k}|^{2}\Bigg|_{j}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|\geq\kappa}\right)$ (73)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{k=1}^{d}\mathbb{E}_{x_{0}}\left(\sigma^{2}_{t}+\sigma^{2}_{t}.{\kappa{\cdot}\sigma^{2}_{t}}.\frac{{\phi}(\kappa{\cdot}\sigma^{2}_{t})}{1-\Phi((\kappa{\cdot}\sigma^{2}_{t}))}\right)$ (74)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{k=1}^{d}\mathbb{E}_{z}\left(2.\sigma^{2}_{t}\right)$ (75)
$\leq\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right)$ (76)
We get Equation equation 74 from Equation equation 73 by using Lemma 7. We get Equation equation 76 from Equation equation 75 from Assumption 1 and by using the upper bound on the Mill’s ration which implies that $\frac{{\phi}(\kappa)}{1-\Phi(\kappa)}\leq{\kappa}+\frac{1}{\kappa}$. We get Equation equation 76 from Equation equation 75 from Assumption 1, which implies that the second moment of $z$ is bounded.
$\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x|_{0}^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (77)
$=\sum_{j=1}^{d}\mathbb{E}_{x_{k}\sim(u_{t})_{k}}\left(|x_{0}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (78)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{k}\sim(u_{t})_{k}|x_{0}}\left(|x_{0}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (79)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{k}\sim(u_{t})_{k}|x_{0}}\left(|x_{0}|^{2}\Bigg|\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)P\left(\Big|\frac{x_{k}-tz_{k}}{1-t}\Big|_{j}\geq\kappa|x_{0}\right)$ (80)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{j=1}^{d}\mathbb{E}_{x_{0}}|x_{0}|^{2}$ (81)
$\leq\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right)$ (82)
Setting $\kappa=\log\frac{dn}{\delta}$ Plugging Equation equation 76, equation 82 into Equation equation 68. Then we have
$|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|\leq$ $\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right),\quad\forall\theta=\{\theta_{k}^{a},\theta_{k}^{b}\}$ (83)
$\leq$ $\mathcal{O}\left(\frac{\delta}{dn}\right)$ (84)
Now plugging Equation equation 83 into Equation equation 62 we get with probability at least $1-2{\delta}$
$|\mathcal{L}_{k}(\theta_{k}^{b})-\mathcal{L}_{k}(\theta_{k}^{a})|$ $\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (85)
Finally, using the Polyak-Łojasiewicz (PL) condition for $\mathcal{L}_{k}(\theta)$, from Assumption 2, we have from the quadratic growth condition of PL functions the following,
$\|\theta_{k}^{a}-\theta_{k}^{b}\|^{2}\leq\mu\left|\mathcal{L}_{k}(\theta_{k}^{a})-\mathcal{L}_{k}(\theta_{k}^{b})\right|,$ (86)
and applying Lipschitz continuity of the velocity fields with respect to parameter $x$
$\|v^{\theta^{a}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ $\leq L_{t}\cdot\|\theta_{k}^{a}-\theta_{k}^{b}\|^{2}$ (87)
$\leq L_{t}\cdot\mu\left|\mathcal{L}_{k}(\theta_{k}^{a})-\mathcal{L}_{k}(\theta_{k}^{b})\right|$ (88)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (89)
Here $L_{t}$ is the Lipschitz parameter of the neural networks. It is always possible to obtain this Lipschitz constant as the quantity $\|v^{a}(x,t)-v^{b}(x,t)\|^{2}\leq L_{t}$ is non-zero only over a finite domain of $x$. Taking expectation with respect to $x$, we obtain the following.
$\mathbb{E}_{x\sim u_{t_{k}}}\mathbb{\|}s^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ $\leq 2\mathbb{E}_{x\sim u_{t}}\mathbb{\|}s^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{b}_{k}}(x,t_{k})-v^{a}_{t}(x)-v^{b}_{t}(x)\|^{2}$
$+2\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{\theta^{a}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ (90)
$\leq 4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{a}_{k}}(x,t_{k})\mathbb{\|}^{2}$ (91)
$+4{\mathbb{E}_{x\sim u_{t}}}\mathbb{\|}s^{\theta^{b}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$
$+4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{a}_{t}(x)-v^{b}_{t}(x)\|^{2}$ (92)
$\leq\mathcal{O}(\kappa^{-2})+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (93)
$\leq\mathcal{O}(\frac{\delta}{dn})+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (94)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (95)
This completes the proof. Note that the quantities $4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}u^{a}_{t}(x)-v^{a}_{t}(x)\mathbb{\|}^{2}$ and $4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}u^{a}_{t}(x)-v^{a}_{t}(x)\mathbb{\|}^{2}$ are bounded in the same manner as is done in Equation equation 82.
C.2 Bounding Optimization Error
The optimization error ($\mathcal{E}_{\text{opt}}$) accounts for the fact that gradient-based optimization does not necessarily find the optimal parameters due to limited steps, local minima, or suboptimal learning rates. This can be bounded as follows.
Proof.
Let $\mathcal{E}_{k}^{\text{opt}}$ denote the optimization error incurred when performing stochastic gradient descent (SGD), with the empirical loss defined by
$\widehat{\mathcal{L}}^{i}_{k}(\theta)=\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (96)
The corresponding population loss is
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{k}}}\left[\left\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}\right],$ (97)
Thus, $\mathcal{E}_{t}^{\text{opt}}$ captures the error incurred during the stochastic optimization at each fixed time step $t$. We now derive upper bounds on this error.
Consider stochastic gradient descent (SGD) iterates indexed by $i=1,2,\dots,n$:
$\theta_{i+1}\;=\;\theta_{i}-\eta_{i}\,\nabla\hat{\mathcal{L}}(\theta_{i}),$
with diminishing stepsizes
$\eta_{i}=\frac{\alpha}{i+\gamma},\qquad\alpha>0,\ \gamma>0,$ (98)
chosen so that
$\alpha\mu>1\quad\text{and}\quad\eta_{i}\leq\frac{1}{L}\ \ \text{for all }i\;\;(\text{e.g.\ take }\gamma\geq\alpha L).$ (99)
Define the expected suboptimality $e_{i}:=\mathbb{E}\bigl[\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr]$.
By Assumption 4 for $\mathcal{L}_{k}$ gives, for $y=\theta_{i+1}=\theta_{i}-\eta_{i}\nabla\hat{\mathcal{L}_{k}}(\theta_{i})$,
$\mathcal{L}_{k}(\theta_{i+1})$ $\leq\mathcal{L}_{k}(\theta_{i})+\Big\langle\nabla\mathcal{L}_{k}(\theta_{i}),\theta_{i+1}-\theta_{i}\Big\rangle+\frac{L}{2}\|\theta_{i+1}-\theta_{i}\|^{2}$
$=\mathcal{L}_{k}(\theta_{i})-\eta_{i}\Big\langle\nabla\mathcal{L}_{k}(\theta_{i}),\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\Big\rangle+\frac{L}{2}\eta_{i}^{2}\bigl\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\bigr\|^{2}.$
Taking conditional expectation given $\theta_{i}$ and using unbiasedness and the variance bound,
$\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]=\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})-\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$
we obtain
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\eta_{i}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\Bigl(\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2}\Bigr).$ (100)
If $\eta_{i}\leq\frac{1}{L}$ then $\eta_{i}+\tfrac{L}{2}\eta_{i}^{2}\leq-\tfrac{\eta_{i}}{2}$, hence
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\frac{\eta_{i}}{2}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$ (101)
By Assumption 2, we have that $\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\geq 2\mu\,(\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star})$, so
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})-\mathcal{L}_{k}^{\star}\mid\theta_{i}\right]\leq\bigl(1-\mu\eta_{i}\bigr)\bigl(\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr)+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$
Taking expectations and substituting $\eta_{i}=\alpha/(i+\gamma)$ yields
$e_{i+1}\;\leq\;\Bigl(1-\frac{\alpha\mu}{i+\gamma}\Bigr)e_{i}\;+\;\frac{\alpha^{2}L\sigma^{2}}{2\,(i+\gamma)^{2}}.$ (102)
Set
$p:=\alpha\mu>1,\qquad b:=\frac{\alpha^{2}L\sigma^{2}}{2}.$
We now prove that any sequence $(e_{i})$ satisfying
$e_{i+1}\leq\Bigl(1-\frac{p}{i+\gamma}\Bigr)e_{i}+\frac{b}{(i+\gamma)^{2}},\qquad p>1,\ \gamma\geq 1,$ (103)
obeys, for all $i\geq 1$,
$e_{i}\;\leq\;\frac{\gamma^{p}e_{1}}{(i+\gamma)^{p}}\;+\;\frac{b}{p-1}\cdot\frac{1}{i+\gamma}.$ (104)
Step 3a: Integrating factor. Define $v_{i}:=(i+\gamma)^{p}e_{i}$. Multiply equation 103 by $(i+1+\gamma)^{p}$:
$v_{i+1}$ $=(i+1+\gamma)^{p}e_{i+1}$
$\leq(i+1+\gamma)^{p}\!\left(1-\frac{p}{i+\gamma}\right)e_{i}\;+\;(i+1+\gamma)^{p}\frac{b}{(i+\gamma)^{2}}.$ (105)
Let $t:=i+\gamma\ (\geq\gamma\geq 1)$. Then
$(i+1+\gamma)^{p}\!\left(1-\frac{p}{i+\gamma}\right)=(t+1)^{p}\!\left(1-\frac{p}{t}\right).$
Step 3b: Bounding the multiplicative factor. By Bernoulli’s inequality, for $u\in[0,1]$, $(1-u)^{p}\geq 1-pu$. With $u=\frac{1}{t}$,
$1-\frac{p}{t}\;\leq\;\left(1-\frac{1}{t}\right)^{p}\;=\;\left(\frac{t-1}{t}\right)^{p}.$
Hence
$(t+1)^{p}\!\left(1-\frac{p}{t}\right)$ $\leq(t+1)^{p}\left(\frac{t-1}{t}\right)^{p}=\left(\frac{t^{2}-1}{t}\right)^{p}\;\leq\;t^{p},$ (106)
since $t^{2}-1\leq t^{2}$. Plugging equation 106 into equation 105 gives
$v_{i+1}\;\leq\;t^{p}e_{i}\;+\;(t+1)^{p}\frac{b}{t^{2}}\;=\;v_{i}\;+\;(t+1)^{p}\frac{b}{t^{2}}.$ (107)
Step 3c: Bounding the additive term. Using the binomial expansion (or the mean-value form), for $t\geq 1$ and $p\geq 1$,
$(t+1)^{p}\leq t^{p}+pt^{p-1}+\frac{p(p-1)}{2}t^{p-2}\;\leq\;t^{p}\!\left(1+\frac{p}{t}+\frac{p(p-1)}{2t^{2}}\right).$
Therefore
$\frac{(t+1)^{p}}{t^{2}}$ $\leq t^{p-2}+p\,t^{p-3}+\frac{p(p-1)}{2}t^{p-4}\;\leq\;\left(1+\frac{p}{\gamma}+\frac{p(p-1)}{2\gamma^{2}}\right)t^{p-2},$ (108)
where in the last inequality we used $t\geq\gamma$ to factor out $t^{p-2}$ and bound the lower powers by constants depending only on $\gamma$ and $p$. Combining equation 107 and equation 108,
$v_{i+1}\;\leq\;v_{i}\;+\;c_{p,\gamma}\,b\,t^{p-2},\qquad c_{p,\gamma}:=1+\frac{p}{\gamma}+\frac{p(p-1)}{2\gamma^{2}}.$ (109)
Sum equation 109 from $j=1$ to i1i-1 (with $t_{j}=j+\gamma$):
$v_{i}\;\leq\;v_{1}\;+\;c_{p,\gamma}\,b\sum_{j=1}^{i-1}(j+\gamma)^{p-2}.$
Since $p>1$, the sum is bounded by an integral:
$\sum_{j=1}^{i-1}(j+\gamma)^{p-2}\;\leq\;\int_{\gamma}^{i+\gamma}t^{p-2}\,dt\;=\;\frac{(i+\gamma)^{p-1}-\gamma^{p-1}}{p-1}\;\leq\;\frac{(i+\gamma)^{p-1}}{p-1}.$
Therefore
$v_{i}\;\leq\;\gamma^{p}e_{1}\;+\;\frac{c_{p,\gamma}\,b}{p-1}\,(i+\gamma)^{p-1}.$
Dividing by $(i+\gamma)^{p}$ yields
$e_{i}\;\leq\;\frac{\gamma^{p}e_{1}}{(i+\gamma)^{p}}\;+\;\frac{c_{p,\gamma}\,b}{p-1}\cdot\frac{1}{i+\gamma}.$ (110)
If one uses the product bound
$\frac{(t+1)^{p}}{t^{p}}\left(1-\frac{p}{t}\right)\;\leq\;\left(1+\frac{p}{t}+\frac{p(p-1)}{2t^{2}}\right)\left(1-\frac{p}{t}\right)\;\leq\;1-\frac{p(p+1)}{2t^{2}},$
valid for $t\geq 1$, then equation 107 improves to $v_{i+1}\leq v_{i}-\frac{p(p+1)}{2t^{2}}v_{i}+b\frac{(t+1)^{p}}{t^{2}}$. Dropping the negative term only helps, and bounding $(t+1)^{p}/t^{2}\leq t^{p-2}$ for $t\geq 1$ gives directly $v_{i+1}\leq v_{i}+b\,t^{p-2}$, leading to equation 104 with the cleaner constant b/(p1)b/(p-1) (i.e., $c_{p,\gamma}=1$) after the same summation step. We will use this sharpened form in the final bound below.
Applying equation 104 to equation 102 with $p=\alpha\mu$ and $b=\frac{\alpha^{2}L\sigma^{2}}{2}$ yields, for all $i\geq 1$,
$e_{i}\;\leq\;\frac{\gamma^{\alpha\mu}\bigl(\mathcal{L}_{k}(\theta_{1})-\mathcal{L}_{k}^{\star}\bigr)}{(i+\gamma)^{\alpha\mu}}\;+\;\frac{\alpha^{2}L\sigma^{2}}{2(\alpha\mu-1)}\cdot\frac{1}{i+\gamma},$
The $O(1/i)$ rate follows since $\alpha.{\mu}\geq 1$.
Thus we obtain
$\mathcal{L}_{k}(\theta_{n})-\mathcal{L}_{k}^{*}\leq\mathcal{O}\left(\frac{1}{n}\right)$ (111)
Thus applying the quadratic growth inequality.
[…]
$\leq\mathcal{O}\left(\frac{1}{n}\right)$ (113)
From lemma 2 we have with probability $1-\delta$ that
$\|s^{a}_{t_{k}}(x,t_{k})-s^{b}_{t_{k}}(x,t_{k})\|^{2}$ $\leq L.\|\theta^{a}_{t}-\theta^{b}_{t}\|^{2}$ (114)
$\leq L.\mu\left|\mathcal{L}_{k}(\theta^{a}_{t})-\mathcal{L}_{k}(\theta^{b}_{t})\right|$ (115)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (116)
Thus we have with probability at least $1-\delta$
[…]
$\leq\mathcal{O}\left({\log}\left(\frac{1}{n}\right)\right)+\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (118)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (119)
Taking expectation with respect to $x\sim p_{t_{k}}$ on both sides completes the proof.
Appendix D Intermediate Lemmas
Lemma 4 (TV bound via Girsanov for reverse diffusions).
Let $X$ and $\tilde{X}$ on $[0,T]$ solve
$dX_{t}=\big(f(X_{t},t)-\sigma^{2}(t)s_{\star}(X_{t},t)\big)dt+\sigma(t)\,d\bar{W}_{t},\qquad d\tilde{X}_{t}=\big(f(\tilde{X}_{t},t)-\sigma^{2}(t)s_{\theta}(\tilde{X}_{t},t)\big)dt+\sigma(t)\,d\bar{W}_{t},$
with the same nondegenerate diffusion $\sigma(t)\in\mathbb{R}^{d\times d}$ (invertible for a.e. $t$) and the same initial law at time $T$. Let $\mathbb{P}$ and $\mathbb{Q}$ be the path measures of $X$ and $\tilde{X}$ on $C([0,T],\mathbb{R}^{d})$. Assume Novikov’s condition
$\mathbb{E}_{\mathbb{Q}}\exp\!\Big(\tfrac{1}{2}\int_{0}^{T}\!\!\|\sigma(t)(s_{\theta}(\tilde{X}_{t},t)-s_{\star}(\tilde{X}_{t},t))\|_{2}^{2}\,dt\Big)<\infty.$
Then
$\mathrm{TV}(\mathbb{P},\mathbb{Q})\ \leq\ \frac{1}{2}\Bigg(\,\mathbb{E}_{\mathbb{Q}}\!\int_{0}^{T}\big\|\sigma(t)\big(s_{\theta}(\tilde{X}_{t},t)-s_{\star}(\tilde{X}_{t},t)\big)\big\|_{2}^{2}\,dt\Bigg)^{\!1/2}.$
Proof.
Write the drift difference as
$\Delta b(x,t)=-\sigma^{2}(t)\big(s_{\star}(x,t)-s_{\theta}(x,t)\big).$
By Girsanov’s theorem (under the stated Novikov condition), $\mathbb{P}\ll\mathbb{Q}$ and the Radon–Nikodym derivative is the exponential martingale driven by $u_{t}=\sigma(t)^{-1}\Delta b(\tilde{X}_{t},t)=\sigma(t)\big(s_{\theta}(\tilde{X}_{t},t)-s_{\star}(\tilde{X}_{t},t)\big).$ The Cameron–Martin formula yields
$\mathrm{KL}(\mathbb{P}\|\mathbb{Q})=\frac{1}{2}\,\mathbb{E}_{\mathbb{Q}}\!\left[\int_{0}^{T}\|u_{t}\|_{2}^{2}\,dt\right]=\frac{1}{2}\,\mathbb{E}_{\mathbb{Q}}\!\left[\int_{0}^{T}\big\|\sigma(t)\big(s_{\theta}(\tilde{X}_{t},t)-s_{\star}(\tilde{X}_{t},t)\big)\big\|_{2}^{2}\,dt\right].$
Applying Pinsker’s inequality $\mathrm{TV}(\mathbb{P},\mathbb{Q})\leq\sqrt{\mathrm{KL}(\mathbb{P}\|\mathbb{Q})/2}$ gives
$\mathrm{TV}(\mathbb{P},\mathbb{Q})\ \leq\ \frac{1}{2}\Bigg(\,\mathbb{E}_{\mathbb{Q}}\!\int_{0}^{T}\big\|\sigma(t)\big(s_{\theta}-s_{\star}\big)\big\|_{2}^{2}\,dt\Bigg)^{\!1/2}.$
Finally, the evaluation map $C([0,T],\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\omega\mapsto\omega(0)$, is measurable, so by data processing for $f$-divergences, $\mathrm{TV}(\mathcal{L}(X_{0}),\mathcal{L}(\tilde{X}_{0}))\leq\mathrm{TV}(\mathbb{P},\mathbb{Q}).$
Let $\{x_{k}\}_{k=0}^{N}$ and $\{\tilde{x}_{k}\}_{k=0}^{N}$ be Euler schemes with the same Gaussian noises,
$x_{k-1}=x_{k}+\big(f_{k}-\sigma_{k}^{2}s_{\star}(x_{k},t_{k})\big)\Delta t_{k}+\sigma_{k}\sqrt{\Delta t_{k}}\,\xi_{k},\quad\tilde{x}_{k-1}=\tilde{x}_{k}+\big(f_{k}-\sigma_{k}^{2}s_{\theta}(\tilde{x}_{k},t_{k})\big)\Delta t_{k}+\sigma_{k}\sqrt{\Delta t_{k}}\,\xi_{k},$
$\xi_{k}\sim\mathcal{N}(0,I)$ i.i.d. Then, with “traj” denoting trajectory measures,
$\mathrm{KL}(\mathrm{traj}_{\star}\|\mathrm{traj}_{\theta})=\frac{1}{2}\sum_{k=1}^{N}\mathbb{E}\Big[\big\|\sigma_{k}\big(s_{\theta}(x_{k},t_{k})-s_{\star}(x_{k},t_{k})\big)\big\|_{2}^{2}\,\Delta t_{k}\Big],$
and hence by Pinsker,
$\mathrm{TV}(\mathrm{traj}_{\star},\mathrm{traj}_{\theta})\ \leq\ \frac{1}{2}\Bigg(\sum_{k=1}^{N}\mathbb{E}\,\big\|\sigma_{k}\big(s_{\theta}-s_{\star}\big)\big\|_{2}^{2}\,\Delta t_{k}\Bigg)^{\!1/2}.$
Lemma 5 (Theorem 26.5 of Shalev-Shwartz & Ben-David (2014)).
Consider data $z\in Z$, the parametrized hypothesis class $h_{\theta},\theta\in\Theta$, and the loss function $\ell(h,z):\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $|\ell(h,z)|\leq c$. We also define the following terms
$L_{D}(h)=\mathbb{E}\ell(h,z)$ (121)
$L_{S}(h)=\frac{1}{m}\sum_{z_{i}\in\mathcal{S}}\ell(h,z_{i})$ (122)
which denote the expected and empirical loss functions respectively.
Then,
With probability of at least $1-\delta$, for all $h\in\mathcal{H}$,
$L_{D}(h)-L_{S}(h)\leq\widehat{R}(\Theta)+4c\sqrt{\frac{2\ln(4/\delta)}{m}}.$ (123)
where $2R(\ell\circ\Theta\circ S)$ denotes the empirical Radamacher complexity over the loss function $\ell$, hypothesis parameter set $\Theta$ and the dataset $\mathcal{S}$
Lemma 6 (Extewnsion of Massart’s Lemma Bousquet et al. (2003)).
Let $\Theta^{{}^{\prime\prime}}$ be a finite function class. Then, for any $\theta\in\Theta^{{}^{\prime\prime}}$, we have
[…]
where $\sigma_{i}$ are i.i.d random variables such that $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=\frac{1}{2}$. We get the second inequality by denoting $L$ as the number of layers in the neural network, $W$ and $B$ a constant such all parameters of the neural network upper bounded by $B$.
Proof.
Let $h_{0}=x$, and for $\ell=0,\dots,L-1$ define the layer recursion
$h_{\ell+1}=\sigma(W_{\ell}h_{\ell}+b_{\ell}),$
where $W_{\ell}\in\mathbb{R}^{n_{\ell+1}\times n_{\ell}}$, $b_{\ell}\in\mathbb{R}^{n_{\ell+1}}$, and $n_{\ell}\leq W$ for hidden layers. We work with the $\ell_{\infty}$ operator norm:
$\|W_{\ell}\|_{\infty}=\max_{i}\sum_{j}|(W_{\ell})_{ij}|\;\leq\;B\,n_{\ell}\;\leq\;BW=\alpha.$
Since $\sigma$ is $1$-Lipschitz with $\sigma(0)=0$, we have $\|\sigma(u)\|_{\infty}\leq\|u\|_{\infty}$ and thus
$\|h_{\ell+1}\|_{\infty}\;\leq\;\|W_{\ell}\|_{\infty}\,\|h_{\ell}\|_{\infty}+\|b_{\ell}\|_{\infty}\;\leq\;\alpha\,\|h_{\ell}\|_{\infty}+B.$
With $\|h_{0}\|_{\infty}\leq d$, iterating this affine recursion yields the standard geometric-series bound
$\|h_{L}\|_{\infty}\;\leq\;\alpha^{L}d\;+\;B\sum_{i=0}^{L-1}\alpha^{\,i}\;=\;\alpha^{L}d\;+\;B\,\frac{\alpha^{L}-1}{\alpha-1}\quad(\alpha\neq 1),$
and for $\alpha=1$, $\|h_{L}\|_{\infty}\leq d+BL$. The scalar output $f(x)$ is either a coordinate of $h_{L}$ or obtained by applying the same $1$-Lipschitz activation to a linear form of $h_{L}$; in either case, $|f(x)|\leq\|h_{L}\|_{\infty}$, giving the stated bound.
For the $\alpha\geq 1$ simplification, use $\sum_{i=0}^{L-1}\alpha^{i}\leq L\alpha^{L-1}$ to obtain
$|f(x)|\;\leq\;\alpha^{L}d+BL\alpha^{L-1}\;=\;(BW)^{L}\!\left(d+\frac{L}{W}\right).$
For $\alpha<1$, since $\alpha^{i}\leq 1$, $\sum_{i=0}^{L-1}\alpha^{i}\leq L$ and hence $|f(x)|\leq d+BL$. Finally, substituting $W=S/L$ gives the size-based form
$|f(x)|\;\leq\;\bigl(BS/L\bigr)^{L}\!\left(d+\frac{L^{2}}{S}\right).$
Lemma 7 (Second Moment of a Symmetrically Truncated Normal).
Let $X\sim\mathcal{N}(\mu,\sigma^{2})$, and let $a>0$. Then the second moment of $X$ conditioned on being outside the symmetric interval $\mu-a,\mu+a]$ is given by
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\mu^{2}+\sigma^{2}+\sigma a\cdot\frac{\phi\left(\frac{a}{\sigma}\right)}{1-\Phi\left(\frac{a}{\sigma}\right)},$
where $\phi(z)=\frac{1}{\sqrt{2\pi}}e^{-z^{2}/2}$ is the standard normal probability density function (PDF), and $\Phi(z)$ is the standard normal cumulative distribution function (CDF).
Proof.
Let $X\sim\mathcal{N}(\mu,\sigma^{2})$. We aim to compute the second moment of $X$ conditioned on the event that it lies outside an interval centered at its mean
$\mathbb{E}[X^{2}\mid|X-\mu|>a]$
This represents the expected squared value of $X$, given that $X$ is in the tails of the distribution (i.e., more than $a$ units away from the mean).
By definition, the conditional expectation is
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\frac{\mathbb{E}[X^{2}\cdot\mathbf{1}_{\{|X-\mu|>a\}}]}{\mathbb{P}(|X-\mu|>a)}$
The numerator integrates $X^{2}$ over the tail regions $(-\infty,\mu-a)\cup(\mu+a,\infty)$, while the denominator is the probability mass in those same regions.
To simplify the integrals, we standardize $X$. Define the standard normal variable
$Z=\frac{X-\mu}{\sigma}\sim\mathcal{N}(0,1)\quad\Rightarrow\quad X=\mu+\sigma Z$
Define $\alpha=\frac{a}{\sigma}$. Then
$|X-\mu|>a\quad\Leftrightarrow\quad|Z|>\alpha$
Our conditional second moment becomes
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\mathbb{E}[(\mu+\sigma Z)^{2}\mid|Z|>\alpha]$
Expanding the square inside the expectation
$(\mu+\sigma Z)^{2}=\mu^{2}+2\mu\sigma Z+\sigma^{2}Z^{2}$
Taking the conditional expectation
$\mathbb{E}[(\mu+\sigma Z)^{2}\mid|Z|>\alpha]=\mu^{2}+2\mu\sigma\mathbb{E}[Z\mid|Z|>\alpha]+\sigma^{2}\mathbb{E}[Z^{2}\mid|Z|>\alpha]$
Since the standard normal distribution is symmetric and the region $|Z|>\alpha$ is also symmetric, we have
$\mathbb{E}[Z\mid|Z|>\alpha]=0$
Thus, the expression simplifies to
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\mu^{2}+\sigma^{2}\mathbb{E}[Z^{2}\mid|Z|>\alpha]$
By definition
$\mathbb{E}[Z^{2}\mid|Z|>\alpha]=\frac{\int_{|z|>\alpha}z^{2}\phi(z)\,dz}{\mathbb{P}(|Z|>\alpha)}=\frac{2\int_{\alpha}^{\infty}z^{2}\phi(z)\,dz}{2(1-\Phi(\alpha))}=\frac{\int_{\alpha}^{\infty}z^{2}\phi(z)\,dz}{1-\Phi(\alpha)}$
Using Intergration by Parts we get,
$\int_{\alpha}^{\infty}z^{2}\phi(z)\,dz=\phi(\alpha)\alpha+1-\Phi(\alpha)$
Therefore
$\mathbb{E}[Z^{2}\mid|Z|>\alpha]=\frac{\phi(\alpha)\alpha+1-\Phi(\alpha)}{1-\Phi(\alpha)}=1+\frac{\alpha\phi(\alpha)}{1-\Phi(\alpha)}$
Substitute back into the expression for $\mathbb{E}[X^{2}\mid|X-\mu|>a]$
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\mu^{2}+\sigma^{2}\left(1+\frac{\alpha\phi(\alpha)}{1-\Phi(\alpha)}\right)$
Recall that $\alpha=\frac{a}{\sigma}$, so the final expression becomes
$\mathbb{E}[X^{2}\mid|X-\mu|>a]=\mu^{2}+\sigma^{2}+\sigma a\cdot\frac{\phi\left(\frac{a}{\sigma}\right)}{1-\Phi\left(\frac{a}{\sigma}\right)}$
Lemma 8 (Linear Growth of Finite Neural Networks).
Let $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ be the output of a feedforward neural network with a finite number of layers and parameters and $\theta\in\Theta$ where $\Theta$ has a finite number of elements. Suppose that each activation function $\sigma:\mathbb{R}\to\mathbb{R}$ satisfies the growth condition
$|\sigma(z)|\leq A+B|z|,\quad\text{for all }z\in\mathbb{R},$
for constants $A,B\geq 0$. Then there exists a constant $C_{\Theta}>0$ such that for all $x\in\mathbb{R}^{d}$,
$|f(x)|\leq C_{\Theta}(1+\|x\|).$
Proof.
We proceed by induction on the number of layers in the network.
Base case: One-layer network.
Let the network be a single-layer function
$f(x)=\sum_{i=1}^{k}a_{i}\,\sigma(w_{i}^{\top}x+b_{i}),$
where $w_{i}\in\mathbb{R}^{d}$, $b_{i}\in\mathbb{R}$, and $a_{i}\in\mathbb{R}$. Then
$|f(x)|\leq\sum_{i=1}^{k}|a_{i}|\cdot|\sigma(w_{i}^{\top}x+b_{i})|.$
Using the growth condition on $\sigma$, we get
$|\sigma(w_{i}^{\top}x+b_{i})|\leq A+B|w_{i}^{\top}x+b_{i}|\leq A+B(\|w_{i}\|\|x\|+|b_{i}|).$
Hence
$|f(x)|\leq\sum_{i=1}^{k}|a_{i}|\left(A+B(\|w_{i}\|\|x\|+|b_{i}|)\right)=C_{0}+C_{1}\|x\|,$
where $C_{0},C_{1}$ are constants depending only on the network parameters. Therefore
$|f(x)|\leq C(1+\|x\|)\quad\text{with }C=\max\{C_{0},C_{1}\}.$
Inductive step.
Assume the result holds for all networks with $L$ layers, i.e., for any such network $f_{L}(x)$,
$|f_{L}(x)|\leq C_{L}(1+\|x\|).$
Now consider a network with $L+1$ layers, defined by
$f_{L+1}(x)=\sum_{j=1}^{k}a_{j}\,\sigma(f_{L}^{(j)}(x)),$
where each $f_{L}^{(j)}(x)$ is an output of a depth-$L$ subnetwork. By the inductive hypothesis
$|f_{L}^{(j)}(x)|\leq C_{j}(1+\|x\|).$
Applying the activation bound
$|\sigma(f_{L}^{(j)}(x))|\leq A+B|f_{L}^{(j)}(x)|\leq A+BC_{j}(1+\|x\|).$
Then
$|f_{L+1}(x)|\leq\sum_{j=1}^{k}|a_{j}|\cdot|\sigma(f_{L}^{(j)}(x))|\leq\sum_{j=1}^{k}|a_{j}|(A+BC_{j}(1+\|x\|))=C_{L+1}(1+\|x\|),$
for some constant $C_{L+1}>0$. This completes the induction.
Examples of Valid Activation Functions
The condition $|\sigma(z)|\leq A+B|z|$ holds for most common activations
  • ReLU: $\sigma(z)=\max(0,z)\Rightarrow|\sigma(z)|\leq|z|$
  • Leaky ReLU: bounded by linear function of $|z|$
  • Tanh: bounded by 1 $\Rightarrow A=1,B=0$
  • Sigmoid: bounded by 1